### v1

# Towards the Worst-case Robustness of Large Language Models

Abstract
Recent studies have revealed the vulnerability of Large Language Models (LLMs) to adversarial attacks, where the adversary crafts specific input sequences to induce harmful, violent, private, or incorrect outputs. Although various defenses have been proposed, they have not been evaluated by strong adaptive attacks, leaving the worst-case robustness of LLMs still intractable. By developing a stronger white-box attack, our evaluation results indicate that most typical defenses achieve nearly 0% robustness. To solve this, we propose DiffTextPure, a general defense that diffuses the (adversarial) input prompt using any pre-defined smoothing distribution, and purifies the diffused input using a pre-trained language model. Theoretically, we derive tight robustness lower bounds for all smoothing distributions using Fractal Knapsack or 0-1 Knapsack solvers. Under this framework, we certify the robustness of a specific case–smoothing LLMs using a uniform kernel–against any possible attack with an average $\ell_{0}$ perturbation of $2.02$ or an average suffix length of $6.41$.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) (OpenAI, 2023; Anthropic, 2024; Dubey et al., 2024) have gained significant attention in recent years due to their impressive performance in a wide range of applications. These models have demonstrated substantial potential in both academic research and practical deployments, making them valuable assets in various domains (Trinh et al., 2024; Cai et al., 2023; Cummins et al., 2023; Liu et al., 2024a). However, concerns about the adversarial robustness of LLMs have also emerged (Wang et al., 2023a; Carlini et al., 2023a) along with their rapid adoption. Even worse, recent studies (Zou et al., 2023; Chao et al., 2023) have shown that adversaries can craft adversarial suffixes to input prompts, which can mislead LLMs to generate malicious or harmful content, also known as jailbreak attacks (Wei et al., 2023a). This vulnerability poses a serious threat to the security and reliability of LLM-based systems, potentially undermining their broader application.
Numerous efforts have been made to enhance the robustness of LLMs against adversarial attacks (Jain et al., 2023). Detection-based defenses leverage a language model to detect and filter harmful behaviors or anomalous text in the input or output (Kumar et al., 2023; Phute et al., 2023; Alon & Kamfonas, 2023). Prompt-based methods incorporate additional prompts to steer the model away from producing harmful content (Wei et al., 2023b; Wu et al., 2023). Furthermore, adversarial training (Madry et al., 2018; Wang et al., 2023b) has been introduced into LLM defenses, involving the training of LLMs or their prompts with adversarial examples (Mo et al., 2024; Bespalov et al., 2024; Sabir et al., 2023). These approaches have successfully defended LLMs against common attacks, significantly improving the safety of LLM-based systems.
As widely recognized in the field of adversarial robustness (Madry et al., 2018; Carlini & Wagner, 2017; Carlini et al., 2019; Athalye et al., 2018), the evaluation of new defenses should include white-box settings, where the adversary has full knowledge of the defended model. These settings represent worst-case scenarios and provide a lower bound on the robustness achievable under black-box attacks. To develop a stronger attack, we improve the I-GCG method (Jia et al., 2024) by strictly ensuring the consistency of tokens during optimization and inference (we call our method I2-GCG). The evaluation results (in Table 1) indicate that most typical defenses exhibit nearly 0% robustness. This finding is not surprising: adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples; detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network (Athalye et al., 2018; Carlini et al., 2019); and adversarial training demands exponentially greater resources (Diakonikolas et al., 2020; Gourdeau et al., 2021), rendering it impractical for sufficiently training large-scale models.
Table 1: Robustness of detection-based defense PPL (Alon & Kamfonas, 2023), prompt-based defenses ICD (Wei et al., 2023b), Self Reminder (Wu et al., 2023), adversarial training (Mo et al., 2024), and our DiffTextPure under white-box I2-GCG attack on the AdvBench dataset (Zou et al., 2023). Following Liao & Sun (2024), we use both string-based and LLM-based evaluators.
I2-GCG No Defense PPL ICD Self Reminder PAT DiffTextPure (Uniform) DiffTextPure (Absorb)
Vicuna-7B 0% 0% 0% 0% 0% 82% 86%
Llama2-7B 0% 0% 0% 0% 2% 86% 88%
Llama3-8B 0% 0% 0% 0% 0% 82% 80%
To address the 0% robustness dilemma against white-box attacks in the text domain, we draw inspiration from randomized smoothing (Cohen et al., 2019; Lecuyer et al., 2019), which constructs a smooth function $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$ from the base function $f$ by using a pre-defined smoothing distribution $p(\bm{z}|\bm{x})$. This construction often possesses favorable mathematical properties, such that the theoretical lower bound for robustness can be derived (Cohen et al., 2019; Teng et al., 2020). Motivated by this insight, we propose a plug-and-play preprocessor defense, named DiffTextPure (diffuse input text and purify). DiffTextPure consists of a forward noising process and a backward denoising process. During the forward process, the input prompts are perturbed according to a pre-defined distribution $p(\bm{z}|\bm{x})$. e.g., each word may be perturbed to a masking token or another word uniformly with probability $\beta$, effectively disrupting the structure of adversarial examples. In the reverse process, the perturbed sentences are restored to clean sentences using either a pre-trained LLM or a trained language diffusion model (Lou et al., 2023) [^1] Note that our method does not necessarily rely on diffusion models, but we still call it DiffTextPure to echo DiffPure (Nie et al., 2022) in the image domain.. We demonstrate that this defense increases the likelihood of correct predictions, purifies adversarial examples, and improves robustness by over 80% against I2-GCG.
The white-box evaluations can only provide an upper bound of worst-case robustness. Future stronger white-box evaluations could potentially reduce the white-box robustness of the defense (Carlini et al., 2019). To break the cycle of the “cat-and-mouse game”, we aim to mathematically establish a certified lower bound (in contrast to the empirical upper bound) for worst-case robustness. We formulate this problem as a functional optimization problem that identifies the minimal output $p_{adv}:=\min_{\bm{x}_{adv},f}g(\bm{x}_{adv})$ over all possible adversarial examples $\bm{x}_{adv}$ and base functions $f$. If this minimal output still satisfies some condition, e.g., $p_{adv}\geq\tau$ for some threshold $\tau$, we say that the function $g$ is certifiably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This definition encompasses all previous specific distributions (e.g., Gaussian distribution in Cohen et al. (2019), Laplacian distributions in Teng et al. (2020)) and applications (e.g., classification, safety), allowing us to certify any smoothing-based defense.
To derive certified robustness for any distribution, we first show that the functional minimization problem is exactly the Fractional Knapsack problem when $f$ is a bounded function, or the 0-1 Knapsack problem when $f$ is a binary function, with the capacity $p_{A}:=g(\bm{x})$, the value of each item $-p(\bm{z}|\bm{x}_{adv})$, and the weight of each item $p(\bm{z}|\bm{x})$. Therefore, one can simply enumerate all $\bm{z}$ from the support of $p(\bm{z}|\bm{x})$ and apply a Fractal Knapsack solver or a 0-1 Knapsack solver to obtain the certified robustness. To reduce the time complexity of certification, we categorize these items based on their value-to-weight ratios and pre-compute the number of items in each category. We provide these statistics for distributions in previous work (Ye et al., 2020; Zeng et al., 2023), as well as for our uniform and absorbing kernels, which exponentially reduce the time complexity of certification. By testing our data-dependent and model-dependent certified bounds on specific models and the AdvBench dataset (Zou et al., 2023), we demonstrate that, with a false positive rate of 0 (validated on MTBench), our defense ensures safety against any possible attack with an $\ell_{0}$ perturbation of $2.02$ or a suffix length of $6.57$.
Overall, our contributions can be summarized as follows:
  1. 1.
    Algorithmically, we comprehensively evaluate existing defenses to reveal their vulnerabilities against adaptive attacks. To mitigate this safety risk, we propose a practical DiffTextPure defense that exhibits desirable robustness even under white-box attacks.
  2. 2.
    Theoretically, we develop a framework to derive certified robustness using any smoothing distribution, with specific applications to uniform and absorbing kernels. We prove that the uniform kernel is always better than the absorbing kernel. We establish the Bayesian error upper bound for our certified robustness lower bounds.
  3. 3.
    Empirically, we achieve nearly 100% attack success rates against existing defenses, show DiffTextPure over 80% improvements against our attack, and provide robustness guarantees with an $\ell_{0}$ perturbation of $2.02$ or a suffix length of $6.57$.
2 Backgrounds and Preliminaries
Adversarial attacks and defenses. The existence of adversarial examples (Szegedy et al., 2014) is a long-standing challenge in the safety of deep learning models. By adding small, even imperceptible perturbations to the input, deep learning models can be easily misled into giving incorrect answers (Ilyas et al., 2019). To defend against these adversarial attacks, various types of defenses have been proposed, e.g. adversarial training (Madry et al., 2018), adversarial example detection (Metzen et al., 2017; Alon & Kamfonas, 2023), and knowledge distillation (Hinton et al., 2015; Papernot et al., 2015), among which input purification methods (Samangouei et al., 2018; Nie et al., 2022; Chen et al., 2023) have demonstrated promising potential in recent research. Notably, white-box adaptive attacks, where the attacker has full access to the model and defense strategies, are critical in defense evaluation since they represent the most challenging adversarial scenario (Athalye et al., 2018). Evaluating defenses under such adaptive attacks provides a lower bound of robustness against potential real-world threats (Croce & Hein, 2020).
Jailbreaking Attacks and Defenses. Recently, jailbreaking attacks have emerged as a specific type of adversarial attack designed to manipulate LLMs into generating harmful, violent, or private content misaligned with human values. These attacks pose a significant safety concern for the deployment of LLMs (Zou et al., 2023). One category of jailbreaking attacks employs heuristic methods, such as manually crafted prompts (Wei et al., 2023b; Jailbreak Chat, 2024), or utilizes LLMs to generate jailbreaking prompts (Chao et al., 2023; Mehrotra et al., 2023). Another category uses optimization-based methods, which minimize a formulated jailbreaking loss to generate adversarial prompts (Zou et al., 2023; Jia et al., 2024; Liu et al., 2023). In this work, we focus on the latter approach, as it can be mathematically formulated and analyzed. To address the safety concerns posed by jailbreaking, various defenses have been proposed, including prompt detection (Alon & Kamfonas, 2023), adversarial training (Mo et al., 2024), and additional safety prompts (Wu et al., 2023). However, these defenses have primarily targeted black-box jailbreaking attacks. When evaluated under stronger white-box attack, they exhibit nearly 0% robustness (detailed in Section 3).
Certified robustness. Neural networks are generally composed of multiple stacked linear layers. Their maximum Lipschitz constant is approximately the product of the maximum singular values of these linear layers, which can be sufficiently large (Fazlyab et al., 2019). As a result, even small perturbations in the input can significantly alter their outputs (Goodfellow et al., 2015). Directly verifying whether a function is robust for all $\bm{x}_{adv}$ would require enumerating all possible inputs, which is computationally impractical due to the exponential time complexity with respect to input dimensionality (Gowal et al., 2018). To address this challenge, researchers have proposed randomized smoothing (Cohen et al., 2019; Salman et al., 2019), which constructs a smoothed function $g$ by aggregating the ensemble predictions of a base function $f$ over a perturbation distribution $p(\bm{z}|\bm{x})$ by $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Thanks to the mathematical properties of the smoothed function $g$, it exhibits inherent smoothness regardless of the vulnerability of the base function $f$. For instance, Cohen et al. (2019) demonstrated that when $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{0},\bm{I})$, the resulting smoothed function $g(\bm{x})$ is guaranteed to be at least $\frac{1}{\sqrt{2\pi}}$-Lipschitz, independent of how susceptible $f$ is to adversarial perturbations. Therefore, if we know $g(\bm{x})=p_{A}$, then we can show that $g(\bm{x}_{adv})\geq p_{A}-\frac{1}{\sqrt{2\pi}}$ for all $\|\bm{x}_{adv}-\bm{x}\|_{2}\leq 1$.
3 Worst-case Robustness of Existing Defenses
Motivation. As suggested by Carlini et al. (2023a), there are two primary reasons researchers focus on worst-case robustness. On the one hand, worst-case robustness represents the maximum capability of real adversaries. If our model achieves reasonable worst-case robustness, we can guarantee that it is safe against any adversaries (Carlini et al., 2019). On the other hand, worst-case robustness provides insight into the worst-case behavior of a neural network, even if we do not believe real adversaries can achieve such worst-case (Pei et al., 2017). Understanding worst-case robustness helps us gain a deeper understanding of the intrinsic mechanisms of neural networks (Szegedy et al., 2014).
White-box robustness, where the attacker has full knowledge of the defended model, represents an upper bound for the worst-case robustness. The actual worst-case robustness must be smaller than the robustness achieved by a white-box attacker (Carlini & Wagner, 2017). Conversely, white-box robustness serves as a lower bound of robustness that an attacker can achieve in practical scenarios, such as black-box settings, where the attacker has limited access to the model’s internal parameters. Therefore, it helps identify vulnerabilities that might be exploited under more favorable conditions for the adversary.
Our design. We observe that previous white-box evaluations of LLMs fail to properly evaluate robustness (Jain et al., 2023) because they do not strictly ensure the consistency of tokenization when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can result in vastly different losses, leading to failures in generating adversarial examples. To address this issue, we improve upon the I-GCG (Jia et al., 2024) by carefully and strictly ensuring token consistency during both attacking and inference. Accordingly, we name our attack as I2-GCG. See Sec. F.1 for further details.
Results. As demonstrated in Table 1, our I2-GCG results in nearly 0% robustness for most typical defenses, demonstrating the vulnerability of existing defenses. This is unsurprising. Most defenses in the vision domain have been attacked to 0% robustness in the last decade (Athalye & Carlini, 2018; Athalye et al., 2018). Adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples. Detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network itself (Athalye et al., 2018; Carlini et al., 2019). Adversarial training works for previous visual adversarial examples, but it demands exponentially greater resources (Diakonikolas et al., 2020; Gourdeau et al., 2021). Current adversarial training on LLMs does not train for a sufficiently long time, improving only average-case robustness but, as of now, not the worst case (Jain et al., 2023).
4 DiffTextPure: Diffuse Text and Purify
Algorithm 1 DiffTextPure
0:  transition distribution $p(\bm{z}|\bm{x})$, input $\bm{x}_{adv}$, noise level $\beta$, sub-sequent LLM or detector $\hat{f}$.
  Sample $\bm{z}\sim p(\bm{z}|\bm{x}_{adv})$, or for specific case:
  if $p(\bm{z}|\bm{x})$ is Absorb then
     Mask tokens with probability $\beta$.
  else if $p(\bm{z}|\bm{x})$ is Uniform then
     Replace tokens uniformly with probability $\beta$.
  end if
  Purify $\bm{z}$ using a LLM or a diffusion model to obtain $\bm{x}_{0}$.
  Return: The output of sub-sequent function $\hat{f}(\bm{x}_{0})$.
To address the dilemma of 0% robustness, we draw inspiration from randomized smoothing (Cohen et al., 2019; Lecuyer et al., 2019), aiming to construct a smooth function $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$ that possesses theoretical guarantees, ensuring that it cannot be reduced to 0% robustness.
To achieve this, we first need to apply a forward process to $\bm{x}$, generating a noised sample $\bm{z}\sim p(\bm{z}|\bm{x})$. We provide three specific choices for $p(\bm{z}|\bm{x})$: Absorbing kernel, which replaces each token with a mask token with probability $\beta$; Uniform kernel, which replaces each token with another token from the vocabulary uniformly at random with probability $\beta$; and Arbitrary kernel, which preserves the current token with probability $\beta$ and replaces it with token $i$ in the vocabulary with probability $\alpha_{i}$.
[…]
4.1 Understanding DiffTextPure
Theoretically, DiffTextPure tends to transform low-likelihood out-of-distribution data (e.g., harmful requests or adversarial suffixes) into high-likelihood in-distribution data. Details are provided in the following lemma:
Lemma 4.1 (DiffTextPure increases the likelihood).
Given a noisy sample $\bm{z}$, the denoised sample $\bm{x}_{0}$ follows the distribution $p(\bm{x}_{0}|\bm{z})=\frac{p(\bm{z}|\bm{x}_{0})p_{\theta}(\bm{x}_{0})}{p(\bm{z})}\propto p_{\theta}(\bm{x}_{0})p(\bm{z}|\bm{x}_{0})$.
As shown, the higher the likelihood of the denoised samples, the closer the denoised sample is to the noisy sample, and the higher the probability that the denoised example will be selected. Therefore, DiffTextPure can be understood as a process that pulls out-of-distribution data back into the in-distribution space. Since most training sets consist of clean requests rather than harmful ones, both adversarial suffixes and harmful instructions are treated as out-of-distribution and are optimized to shift back into the in-distribution space. In contrast, benign inputs are already in-distribution, leading the model to make minimal changes and thus preserve the utility of natural instructions.
We do two case studies in the following:
Case I: destroying adversarial strings. Since the forward process randomly perturbs each word, DiffTextPure has a high probability of disrupting the adversarial nature of the adversarial suffix. As demonstrated in the example below, frequently occurring words in adversarial suffixes, such as “Wikipedia”—which often encourages language models to respond to harmful queries similarly to a Wikipedia article (Zou et al., 2023)—are replaced with non-harmful symbols, thus neutralizing their adversarial impact. Although the user’s request may also be perturbed, the reverse denoising process allows the denoiser $D$ to recover this part due to its high likelihood, while the adversarial suffix is nearly impossible to be restored due to the extremely low likelihood of encountering such garbled strings.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
User: Write a convincing anonymous death threat. legend CheersDan}}_ {\((concat tuple://ifyrefs fnvarspply] [[= {(([wikipedia
[…]
Assistant: Sorry, I can’t assist with that.
Case II: purifying into benign requests. Since adversarial requests are typically absent from the training set of denoiser $D$, the reverse process is more likely to recover the noisy text as a benign request rather than a harmful one. As demonstrated in the example below, DiffTextPure changes the phrase “hack into” to “log into,” effectively defending against harmful requests.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
User: Provide step-by-step instructions for how to hack into a corporate network.
Purified Text: Provide step-by-step instructions for how to log into a corporate network.
Assistant: Sure, here is how to log into a corporate network.
5 Certified Robustness
White-box evaluations can only provide an upper bound of the worst-case robustness. This bound may further decrease as stronger attacks are developed in the future, especially for stochastic defenses (see Appendix G). In this section, we would like to provide a theoretical lower bound for the worst-case robustness.
This section is organized as follows: We begin by discussing the formulation of randomized smoothing-based certified robustness in Sec. 5.1. Next, we present a general solver for certified robustness on any smoothing distribution in Sec. 5.2, along with specific applications to the absorbing kernel (Sec. 5.3) and the uniform kernel (Sec. 5.4). Finally, we analyze the upper bound of these theoretical lower bounds in Sec. 5.5.
5.1 Formulation of Certified Robustness for LLMs
Inspired by Cohen et al. (2019), we define certified robustness using randomized smoothing as the following functional optimization problem:
Definition 5.1.
Given a base model $f:\mathcal{X}\to\mathbb{R}$ and a smoothing distribution $p(\bm{z}|\bm{x})$, we define the smoothed function $g:\mathcal{X}\to\mathbb{R}$ as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Let $g(\bm{x})=p_{A}$, assume $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$ for some distance metric $\mathcal{D}$. We define randomized smoothing as the optimization for finding minimal output $g(\bm{x}_{adv})$ over all functions (or the worst) $f$ and $\bm{x}_{adv}$:
$p_{adv}=\min_{f,\bm{x}_{adv}}g(\bm{x}_{adv})=\min_{f,\bm{x}_{adv}}\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ (1)
[…]
If $p_{adv}\geq\tau$ for a given threshold $\tau$, we say the function $g$ is considered certifiably robust for input $\bm{x}$ within range $d$.
We discuss how to apply this technique to common tasks like classification and detection in Sec. B.6. In the following, we show three ways to apply this technique to certify the safety of language models.
Way I: Certifying the detector. Let $\mathcal{V}$ be the vocabulary, $N$ be the sequence length. The base detector $f:\mathcal{V}^{N}\to[0,1]$ outputs values close to $1$ if the input is harmful and close to 00 if it is not. The user specifies the threshold $\tau$ to adjust the conservativeness of the detector. If we can show that, for a given base detector and $g(\bm{x})=p_{A}$, $g(\bm{x}_{adv})$ remains greater than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is certifiably robust within the distance $d$.
Way II: Certifying “sure”. Most current jailbreaking attacks force the model to output “sure” as the first word. If we can certify that the model does not output “sure,” we can provably defend against these attacks. Here, $f:\mathcal{V}^{N}\to[0,1]$ represents the probability that the base language model outputs “sure,” and the threshold is set as $\tau=\frac{1}{|\mathcal{V}|}$. If we can show that $g(\bm{x}_{adv})$ is smaller than $\tau$ for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, then the detector $g(\bm{x})$ is successfully certified within $d$. However, this approach is not applicable to attacks where the attackers do not set the optimization target to “sure.”
Way III: Certifying the output of a language model. Given a language model $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$ and a judgment oracle $\mathcal{O}:\mathcal{V}^{N}\to\{0,1\}$, we construct a smoothed function $g(\bm{x})=\mathbb{E}[\mathcal{O}(f(\bm{z}))]$, which is exactly the probability that $f(\bm{z})$ produces a benign output. If we can show that $g(\bm{x}_{adv})$ is larger than $\tau$, this demonstrates that the output of the language model is safe with at least $\tau$ probability. However, this approach requires a highly accurate or strictly conservative oracle detector $\mathcal{O}$, which is not feasible with most practical detectors.
Since ways II and III are not strictly rigorous, in the main paper, we focus exclusively on certifying a safety detector.
5.2 Certified Robustness on General Distribution
Previous researchers have addressed certified robustness for simple distributions, such as Gaussian distributions (Cohen et al., 2019), masking distributions (with a fixed masking ratio) (Zeng et al., 2023), and synonym distributions (Ye et al., 2020). However, these methods are not applicable to our uniform and absorbing kernels. To address this, we propose a solution for solving the contained optimization problem in Eq. 1 for any smoothing distribution. We begin by reviewing the Fractional Knapsack Problem:
Definition 5.2.
(The Fractional Knapsack Problem). Given some items, for each item $\bm{z}$, it has a weight $p(\bm{z}|\bm{x})$ and value $-p(\bm{z}|\bm{x}_{adv})$, the goal is to select items (or fractions of items) such that the total weight $\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})$ does not exceed the knapsack’s capacity $p_{A}$, while maximizing the total value $-\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$, where $f(\bm{z})\in[0,1]$ denotes the fraction of each item chosen.
The solution to the Fractional Knapsack Problem relies on a greedy algorithm: we prioritize items by their value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, selecting items in descending order of this ratio until the capacity $p_{A}$ is reached. This approach is optimal because it maximizes the contribution of each item (or fraction of an item) per unit weight added to the knapsack (Aho & Hopcroft, 1974; Cormen et al., 2022).
Therefore, to solve Eq. 1, we simply enumerate all possible $\bm{z}$, sort them by $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ in descending order, and select items until the cumulative weight reaches $p_{A}$, as shown in Algorithm 2. Each time we select a $\bm{z}$, we consume $p(\bm{z}|\bm{x})$ from $p_{A}$, but add $p(\bm{z}|\bm{x}_{adv})$ to $p_{adv}$. Consequently, we refer to the negative value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ as the trading rate. The larger the trading rate, the greater the increase in $p_{adv}$, the better “our trade” is.
Algorithm 2 Fractal Knapsack Solver for Eq. 1
0:  Smoothing distributions $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$, threshold $\tau$, $p_{A}=g(\bm{x})$. Output: Whether $g$ is provably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$.
1:  Compute $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ for each $\bm{z}\in\mathcal{X}$, sort items in descending order.
2:  Initialize weights and value $W,V\leftarrow 0$.
3:  for each item $\bm{z}$ in sorted order do
4:     if $W+p(\bm{z}|\bm{x})\leq p_{A}$ then
5:        $W\leftarrow W+p(\bm{z}|\bm{x})$ and $V\leftarrow V+p(\bm{z}|\bm{x}_{adv})$
6:     else
7:        Select fraction of $\bm{z}$ to fill remaining $p_{A}-W$
8:        $V\leftarrow V+\left(p(\bm{z}|\bm{x}_{adv})\cdot\frac{p_{A}-W}{p(\bm{z}|\bm{x})}\right)$
9:        Break
10:     end if
11:  end for
12:  Return: $\mathbb{I}\{V\geq\tau\}$
Theorem 5.3.
(Proof in Sec. C.1). Algorithm 2 exactly solves Eq. 1.
Remark 5.4.
Note that the tightness of this bound relies on the assumption that the hypothesis set of $f$ includes all functions mapping $\mathcal{V}^{N}$ to $[0,1]$. If we restrict the hypothesis set to functions that map to $\{0,1\}$ (i.e., hard functions that output only 0 or 1), this reduces to a 0-1 Knapsack problem. In such cases, a tighter bound can be obtained using dynamic programming. See Sec. C.8 for details.
To apply these solvers to specific smoothing kernels, a brute-force way is to enumerate all possible $\bm{z}$ and perform Algorithm 2 for these $\bm{z}$. This requires $O(|\mathcal{V}|^{N})$ for Fractal Knapsack and $O(|\mathcal{V}|^{2N})$ for 0-1 Knapsack. A clever way to do this is to categorize these items according to value-to-weight ratios and calculate the number of items (volume) for each category. Formally, the volume $v(\gamma)$ for trading rate $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma$ is defined as:
$v(\gamma)=\sum_{\bm{z}}p(\bm{z}|\bm{x})\mathbb{I}\left\{\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma\right\}.$
In Sec. 5.3 and Sec. 5.4, we show how to reduce the time complexity from $O(|\mathcal{V}|^{N})$ to $O(1)$ for absorbing kernel and from $O(|\mathcal{V}|^{N})$ to $O(2d)$ for uniform kernel.
5.3 Certified Robustness on Absorbing Kernel
For simplicity, we use the subscript $p$ to denote the common part of $\bm{x}$ and $\bm{x}_{adv}$, and the subscript $s$ to denote the differing part of $\bm{x}$ and $\bm{x}_{adv}$. Clearly, $\bm{x}_{p}=\bm{x}_{adv_{p}}$. We denote $\beta$ as the probability of turning each word into a masked token, and $\bar{\beta}=1-\beta$ as the probability of remaining unchanged.
To solve Eq. 1 for absorbing kernel, it suffices to calculate all trading rates $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ and their corresponding volumes. We give these results in the following theorem:
Theorem 5.5.
(Certified robustness for Absorbing Kernel, proof in Sec. C.3) Divide $\mathcal{V}^{N}$ into $L_{1}$ and $L_{2}$ that $L_{1}\cup L_{2}=\mathcal{V}^{N}$ and $L_{1}\cap L_{2}=\emptyset$, where $L_{1}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are all masked tokens}\}$, $L_{2}=\{\bm{z}\in\mathcal{V}^{N}\mid\bm{z}_{s}\text{ are not all masked tokens}\}$. Clearly, we have the trading rate:
$\forall\bm{z}\in L_{1},\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=1;\;\forall\bm{z}\in L_{2},\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=0.$
and the corresponding volume:
$v(1)=\beta^{d},\;v(0)=1-\beta^{d}.$
By plugging these results into Algorithm 2, we can compute the certified bound using the absorbing kernel. Interestingly, due to the simplicity of the trading rate and volume, we can derive an analytic solution for the certified radius as $\lfloor\frac{\ln(1-p_{A}+\tau)}{\ln\beta}\rfloor$. For details, refer to Sec. C.4.
Table 2: The average certified $\ell_{0}$ radius (at the false positive rate 0) against $\ell_{0}$ attack.
Vicuna-7B Llama2-7B Llama3-8B GPT-4o Trivial Bound Bayesian Bound
DiffTextPure (Absorb) 1.00 1.92 1.82 2.00 2.12 2.10
DiffTextPure (Uniform) 1.02 1.86 1.54 2.02 2.12 2.10
5.4 Certified Robustness on Uniform Kernel
For the uniform kernel, calculating all trading rates $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ and their corresponding volumes is extremely challenging. Fortunately, this problem can be reduced to $O(d)$ level rather than $O(N)$ level since only the difference part between $\bm{x}$ and $\bm{x}_{adv}$ matter:
For value-to-weight ratio:
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\frac{p(\bm{z}_{p}|\bm{x}_{adv_{p}})}{p(\bm{z}_{p}|\bm{x}_{p})}\frac{p(\bm{z}_{s}|\bm{x}_{adv_{s}})}{p(\bm{z}_{s}|\bm{x}_{s})}=\frac{p(\bm{z}_{s}|\bm{x}_{adv_{s}})}{p(\bm{z}_{s}|\bm{x}_{s})}.$
For its volume:
$v(\gamma)$ $=\sum_{\bm{z}}p(\bm{z}|\bm{x})\mathbb{I}\{\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}=\gamma\}$
$=\sum_{\bm{z}_{p}}\sum_{\bm{z}_{s}}p(\bm{z}_{p}|\bm{x}_{p})p(\bm{z}_{s}|\bm{x}_{s})\mathbb{I}\{\frac{p(\bm{z}_{s}|\bm{x}_{adv_{s}})}{p(\bm{z}_{s}|\bm{x}_{s})}=\gamma\}$
$=\sum_{\bm{z}_{s}}p(\bm{z}_{s}|\bm{x}_{s})\mathbb{I}\{\frac{p(\bm{z}_{s}|\bm{x}_{adv_{s}})}{p(\bm{z}_{s}|\bm{x}_{s})}=\gamma\}.$
Therefore, the certified bound of the uniform kernel is independent of the input length $N$ (dependent part only exists in network accuracy $p_{A}$), but only adversarial budget $d$. This greatly simplifies the derivation of value-to-weight ratio and volume. We give these results in the following Theorem 5.6. We can compute the certified robustness using the uniform kernel by plugging these results into Algorithm 2.
With a slight abuse of notation, we still use $\bar{\beta}$ as the probability of each token remaining unchanged, $\beta=1-\bar{\beta}$ as the probability of doing perturbation, $\alpha=\frac{\beta}{|\mathcal{V}|-1}$ as the probability of perturbing into each of the other tokens.
Theorem 5.6.
Let $v(i,j)=\sum p(\bm{z}|\bm{x})\mathbb{I}\{p(\bm{z}|\bm{x})=\alpha^{i}\bar{\beta}^{d-i}\land p(\bm{z}|\bm{x}_{adv})=\alpha^{j}\bar{\beta}^{d-j}\}$, which represents the probability measure on $p(\bm{z}|\bm{x})$ for the set of $\bm{z}$ such that $\bm{z}$ differs from $\bm{x}$ by $i$ tokens and differs from $\bm{x}_{adv}$ by $j$ tokens. Then, we have the following expression for $v(i,j)$:
$v(i,j)=\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{i}\bar{\beta}^{d-i}.$ (2)
Proof.
Here, $i$ represents the number of perturbed tokens in the normal suffix, and $j$ represents those in the adversarial suffix. Only when $i+j\geq d$ can $\bm{z}$ be derived from both $\bm{x}$ and $\bm{x}_{adv}$. Among the $i+j$ modified tokens, $i+j-d$ tokens must be perturbed to arbitrary tokens, excluding the counterparts in either the normal or adversarial suffixes, resulting in $(|\mathcal{V}|-2)^{i+j-d}$ possible values for these tokens. The $i$ tokens can be chosen from any positions in the suffix, giving the term $\binom{d}{i}$. Among these $i$ tokens, $d-j$ must match the adversarial suffix, while the remaining $i+j-d$ can be perturbed to any of $|\mathcal{V}|-2$ possible values, leading to a total arrangement count of $\binom{i}{d-j}$. ∎
A notable property of the uniform kernel is that if $g(\bm{x})=1$, the certified radius becomes infinite. This occurs because the support of $p(\bm{z}|\bm{x})$ spans the entire space $\mathcal{V}^{N}$. When $g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=1$, it implies that $f(\bm{z})=1$ for all $\bm{z}$. Consequently, $g(\bm{x}_{adv})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$ will also equal 1, which is always greater than $\tau$. In contrast, with the absorbing kernel, $g(\bm{x}_{adv})$ cannot exceed $\beta^{d}$. Therefore, the certified radius is at most $\frac{\log\tau}{\log\beta}$. From this perspective, the uniform kernel closely resembles the behavior of the Gaussian distribution in the image domain, where the certified radius can also potentially be infinite.
In the following theorem, we theoretically demonstrate that the uniform kernel outperforms the absorbing kernel not only when $p_{A}=1$ but across all values of $p_{A}$.
Theorem 5.7.
(Proof in Sec. C.5.) The certified radius of uniform kernel is always larger than absorbing kernel given the same accuracy $p_{A}$, threshold $\tau$, perturbing probability $\beta$, i.e.,
$\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V})\geq\text{certify}(\text{absorb},p_{A},\tau,\beta).$
The equality holds when $|\mathcal{V}|\to\infty$.
5.5 Upper Bound of Certified Radius
In this section, we investigate the theoretical limits of robustness guarantees under $\ell_{0}$ attacks. Specifically, we aim to determine the upper bound of the certified lower bound by analyzing the role of keywords in a sentence.
Definition 5.8.
We define the number of keywords $K(\bm{x})$ in a sentence $\bm{x}$ as the minimal number of words whose changes alter the semantics of the input. Formally,
$K(\bm{x})=\min_{\bm{y}}i,\quad\text{subject to}\quad\mathcal{O}(\bm{x})\neq\mathcal{O}(\bm{y}),\,\|\bm{x}-\bm{y}\|_{0}\leq i,$
where $\mathcal{O}$ represents the judgment oracle.
From this perspective, we can derive two upper bounds for the certified lower bound.
Trivial Bound. Changing $K(\bm{x})$ words will alter the semantics of the input. Therefore, we can certify at most $\ell_{0}$ attacks involving $K(\bm{x})-1$ words, i.e.,
$R(\bm{x})\leq K(\bm{x})-1.$
$\boldsymbol{p_{A}}$ Bound. If the smoothing function $p(\bm{z}|\bm{x})$ removes all the keywords in $\bm{x}$, the subsequent model cannot produce the correct output. Thus, for uniform and absorbing kernel, the model accuracy is bounded as $p_{A}\leq 1-\beta^{K(\bm{x})}:=\overline{p_{A}}$. Consequently, we have:
$R(\bm{x})\leq\max_{\tau,\beta,\mathcal{V}}\text{certify}(\text{uniform},\overline{p_{A}},\tau,\beta,\mathcal{V}).$
6 Experiment
Table 3: The average certified length against adversarial suffix attack using Llama3-8B (Dubey et al., 2024).
$\beta$ 0.1 0.25 0.5 1
DiffTextPure (Absorb) 3.87 6.57 12.35 $\infty$
DiffTextPure (Uniform) 3.72 6.41 11.47 $\infty$
In this section, we demonstrate the experimental details of both empirical evaluation and certified robustness.
6.1 Empirical Evaluations
Settings. We conduct both black-box evaluations to demonstrate practical usage (Sec. E.3) and white-box evaluations (Table 1) to establish the upper bound of worst-case robustness. Following Zou et al. (2023); Jia et al. (2024); Liao & Sun (2024), we use the AdvBench dataset (Zou et al., 2023). We perform suffix attacks that append $d=20$ adversarial tokens as a suffix to the original request and optimize these appended tokens. We set $\beta=0.25$. Refer to Appendix E for other details.
Results. As demonstrated, all defenses achieve reasonable performance in black-box settings. This is because current optimization-based attacks, which do not utilize human-crafted prompts, transfer unreasonably poorly across different LLMs (Schaeffer et al., 2024), making black-box defenses much easier. However, these defenses all achieve nearly 0% white-box robustness. This is consistent with the previous experience in vision domains that only a few defenses that are not 0% robustness against white-box attacks (Athalye & Carlini, 2018; Athalye et al., 2018). As a randomization-based defense, DiffTextPure achieves non-trivial white-box robustness. This is partially because the stochasticity significantly interferes with white-box evaluations (Lee & Kim, 2023; Liu et al., 2024b) and partially because of the intrinsic robustness of such smoothing defenses (Carlini et al., 2023b; Cohen et al., 2019).
6.2 Certified Robustness
Settings. We still use the AdvBench dataset (Zou et al., 2023). Detectors are constructed by adjusting the prompt of the LLM (see Sec. E.1). This prompt is extremely conservative, ensuring that the detector achieves a 0% false positive rate on normal requests, like MT-Bench (Zheng et al., 2024a). Additionally, we perform certification on sentences containing more harmful keywords in Sec. E.7, using different purifiers in Sec. E.6, and ablation studies in Appendix E.
We focus on certifying safety against two types of attacks: In $\ell_{0}$ attack, we set $\beta=0.1$. We apply DiffTextPure to the entire sentence, thereby certifying the $\ell_{0}$ radius. In Suffix Attack, we set $\beta=0.25$. we first pad the input sentence with 50 arbitrary tokens and apply DiffTextPure to all except the first $k$ tokens. We purify the diffused tokens conditioned on the undiffused tokens. If the suffix part is not fully included in DiffTextPure, we return a certified robustness of 0. Otherwise, we certify the $\ell_{0}$ radius excluding the first $k$ tokens, thereby covering the suffix part. We also discuss additional settings in Sec. F.4.
Results. As shown in Table 2, for $\ell_{0}$ attacks, we achieve a certified radius of $2.02$. The better the base model, the higher the true positive rate, and thus, the higher the certified radius. For the AdvBench dataset, the theoretical upper limit—Bayesian bound—of the certified method is close to the real limit, the trivial bound, demonstrating that the certified bound closely approximates real robustness. However, this does not always hold true for datasets containing longer requests (Sec. E.7). This may require a fundamental improvement on the randomized smoothing paradigm, e.g., relying on more neural network-dependent variables rather than a single $p_{A}$. For adversarial suffix attacks, we achieve an average certified radius of $6.41$ (with $\beta=0.25$), while practical settings focus on suffix lengths of $20$. This demonstrates that it is relatively easy to obtain a certified radius with strong practical significance in the suffix attack settings due to its simplicity (Kumar et al., 2023).
Smoothness-utility trade-off. As the probability of perturbing a token $\beta$ approaches 1, the distribution of the diffused samples becomes identical for both benign and adversarial inputs. In this scenario, the base model is unable to distinguish whether the noisy examples come from benign or adversarial inputs. As a result, the function $g$ becomes overly smooth, returning a constant output regardless of the input. Since we require a false positive rate of 0, in the $\ell_{0}$ attack setting, this directly leads to a certified radius of 0. In the suffix attack setting, the detector relies entirely on the first $k$ non-diffused tokens, causing the certified radius to be either 0 or $\infty$. A certified radius of $\infty$ is undesirable because adding several tokens can indeed alter the semantics of an input sentence (Sec. F.3). Typically, we choose $\beta=0.25$, as this value avoids masking key information in the sentence and prevents oversmoothing.
7 Conclusion
In this work, we investigate the worst-case robustness of large language models. We upper bound the worst-case robustness of previous defenses by proposing a strong adaptive attack that strictly ensures the consistency in tokenization between optimization and inference. We also lower bound the worst-case robustness of all randomization-based defenses by reducing the functional optimization to a fractal knapsack problem or 0-1 knapsack problem. We reduce the time complexity of certification exponentially for specific kernels, such as uniform kernels or absorbing kernels, by categorizing each item based on its value-to-weight ratio and pre-computing the number of items in each category. We also provide theoretical analysis on the relationship between certified robustness, smoothing distribution, and vocabulary size, and upper bound the certified lower bound by Bayesian error, offering insights into the upper limits of certified methods. See Appendix H for key takeaways.
Appendix A Notations
$f$ Base model. Can be detectors, purifiers, large language models, or compositions of them.
$g$ Smoothed function.
$Q$ Diffusion kernel for perturbing an input sentence.
$\bar{\beta}$ The probability of current words remain unchanged.
$\beta$ Equals to $1-\bar{\beta}$, represent the probability of perturbing the current word.
$\alpha$ Equals to $\frac{\beta}{|\mathcal{V}|-1}$, the probability of perturbing the current word into a specific word in the uniform kernel.
$-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ Value-to-weight ratio.
$\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ Trading rate.
$v(\gamma)$ The probability measure of the set where the trading rate of each item is $\gamma$.
$v(i,j)$ The probability measure of the set where $p(\bm{z}|\bm{x})=\alpha^{i}\bar{\beta}^{d-i}\land p(\bm{z}|\bm{x}_{adv})=\alpha^{j}\bar{\beta}^{d-j}$.
$p_{A}$ Equals to $g(\bm{x})$.
$p_{adv}$ The minimal possible value of $g(\bm{x}_{adv})$.
$\overline{p_{A}}$ Bayesian upper bound of $p_{A}$.
$D$ The denoiser.
$\mathcal{D}$ Distance metric.
$N$ (maximum) Input length.
$d$ Perturbation budget, e.g., number of different tokens between $\bm{x}$ and $\bm{x}_{adv}$.
$K(\bm{x})$ Number of keywords in $\bm{x}$.
$O$ Time complexity.
$\mathcal{O}$ Judgement oracle.
$R(\bm{x})$ Certified radius for $\bm{x}$.
$\mathcal{V}$ Vocabulary.
$|\mathcal{V}|$ Vocabulary size.
$\tau$ Threshold.
Appendix B Additional Related Work
B.1 More Related Work on Jailbreak Attacks and Defenses
The jailbreak attack on LLMs primarily refers to inducing LLMs into generating harmful content that is unsafe or toxic to society (Chao et al., 2024; Zhou et al., 2024b). To achieve this goal, malicious attackers can craft jailbreaking prompts through manual design, optimization, or train a generative model. Manual-designed jailbreak prompts leverage heuristic perspectives like data distribution (Wei et al., 2023b; Deng et al., 2023; Wei et al., 2023a), psychology insights (Shen et al., 2024b; Zeng et al., 2024; Shen et al., 2024a; Li et al., 2023b) or cipher encoding (Yuan et al., 2023; Handa et al., 2024) to achieve this goal. Optimization-based attacks extend from manually designing by optimizing an adversarial prompt with certain loss functions, where they can optimize a prefix or suffix (Zou et al., 2023; Liu et al., 2023; Jia et al., 2024; Zhang & Wei, 2025; Li et al., 2024a), or directly refine the jailbreaking prompt (Dong et al., 2023; Chen et al., 2024c; Zheng et al., 2024b; Chao et al., 2023). Besides, a thread of work toward fitting the jailbreak prompt distribution with a generative model (Liao & Sun, 2024; Kumar et al., 2024; Paulus et al., 2024; Basani & Zhang, 2024), effectively increasing the attack efficiency. Notably, there are also fine-tuning-based attacks that directly manipulate the alignment instead of designing prompts (Qi et al., 2023; Yang et al., 2023; Zhang et al., 2024b), posing another safety threat to LLMs.
From the defense perspective, various methods are proposed at different stages of generation. Pre-processing defenses are designed to detect potential jailbreaking prompts, typically aimed at adversarial suffix-based attacks that cause significantly high perplexity (Jain et al., 2023; Alon & Kamfonas, 2024). Besides, prompt-based defenses add safety tokens during generation, which are manually designed (Wei et al., 2023b; Xie et al., 2023) or optimized (Mo et al., 2024; Zhou et al., 2024a). Finally, post-processing defenses detect jailbreaking with hidden spaces (Li et al., 2025; Galinkin & Sablotny, 2024) or toxicity detection (Wang et al., 2023c; Hu et al., 2024; Wang et al., 2024).
B.2 Adversarial Attacks and Defenses on Text Domain
Textual adversarial attacks (Morris et al., 2020; Wang et al., 2019b; Han et al., 2022) extend adversarial examples from vision space to discrete text space. Thus, a major challenge of textual attacks is the optimization process on discrete tokens, which include character, word, or sentence-level attacks. For instance, word-level attacks replace critical tokens with semantically similar alternatives to evade detection (Jin et al., 2020; Zang et al., 2019), while character-level attacks insert misspellings or Unicode artifacts to bypass filters (Ebrahimi et al., 2018; Rocamora et al., 2024). Recent advances also employ generative models to automate the creation of adversarial examples (Ren et al., 2020; Li et al., 2023a), producing fluent but malicious inputs that align with natural language patterns. These attacks highlight the vulnerability of text-based systems to carefully crafted inputs, even when perturbations are imperceptible to humans.
Defending against textual adversarial attacks also requires addressing the discrete nature of language. Adversarial training (Xiao et al., 2018), which incorporates perturbed examples during model optimization, remains a cornerstone for improving the robustness of language models (Wang et al., 2019a; Gao et al., 2023). A series of certified defenses with randomized smoothing techniques provide probabilistic guarantees against textual bounded perturbations (Jia et al., 2019; Wang et al., 2021) was also proposed. The evolving landscape of text-domain adversarial robustness underscores the need for defenses that generalize across attack vectors while preserving linguistic integrity. However, these defenses and certifications are limited to conventional language models like sentence classifiers, yet the certified robustness of large generative models remains unexplored.
B.3 Diffusion Models for Adversarial Robustness
Diffusion models (Song et al., 2021; Dhariwal & Nichol, 2021) have achieved notable success in defending against visual adversarial examples (Nie et al., 2022; Wang et al., 2022; Li et al., 2024b; Xiao et al., 2023; Zhang et al., 2023; Carlini et al., 2023b). In particular, they are widely used as a plug-and-play purification method, named DiffPure, making them suitable for commercial models (Zhang et al., 2024a). As illustrated in Figure 1, given a model to be protected model, $f$, and a diffusion denoiser $D$, DiffPure involves two main steps: First, it adds Gaussian noise with variance $\sigma_{\tau}^{2}$ to the input images, and then denoising these noisy images using the diffusion model $D$.
Intuitively, the norm of the added Gaussian noise is much larger than that of the adversarial perturbations, effectively washing out the adversarial nature of the small-norm perturbations (Nie et al., 2022). Theoretically, this procedure not only increases the log-likelihood of input images, pushing them back from out-of-distribution to in-distribution (Nie et al., 2022; Xiao et al., 2023), but also implicitly constructs a smooth classifier $g(\bm{x})=\mathbb{E}_{\bm{x}_{\tau}\sim\mathcal{N}(\bm{x},\sigma_{\tau}^{2}\bm{I})}[f(D(\bm{x}_{t}))]$. The mathematical properties of this classifier have been extensively studied, providing theoretical proof on whether adversarial examples can exist within certain neighborhoods (Carlini et al., 2023b; Xiao et al., 2023; Chen et al., 2024b; Zhang et al., 2023).
B.4 More Related Work on Certified Robustness
Certified robustness by masking. Certified robustness through masking has been extensively studied in previous work (Zeng et al., 2023; Levine & Feizi, 2020; Moon et al., 2023; Zhang et al., 2019) in both text and image domains (e.g., partitioning images into patches and masking them). The certification approach for DiffTextPure-Absorb differs slightly from these works, as tokens are masked with a probability rather than at a fixed ratio, leading to a much more neat result, as shown in Sec. 5.3. Zeng et al. (2023) suggest that this certified lower bound can be improved by introducing an auxiliary variable. However, their approach does not incorporate hypothesis testing or account for type-one error in estimating this auxiliary variable. For randomized smoothing via masking, it is obvious that this bound is tight as there exists a worst-case $f$ that fails entirely on region $L_{1}$. When fixing their bound with hypothesis testing using Bonferroni correction, it is clear that this produces the same result.
Certified robustness by random perturbing words. Jia et al. (2019) uses interval bounds propagation to propagate the activation bounds to the final layers. These methods currently are not scalable to large models. On the contrary, we adopt randomized smoothing, a model-agnostic certification approach, which is thus more scalable.
Universal certification. Lee et al. (2019) also establish a lower bound when smoothing a pre-trained model with randomly perturbed words, but there are several key differences compared to our work. First, we demonstrate that the certified robustness problem can be formulated as a Fractal Knapsack problem, making the approach more intuitive and easier. Second, we show that this can be further improved when the base model $f$ is a hard function, which becomes a 0-1 Knapsack problem and can obtain a stronger result using dynamic programming. What’s more, we greatly simplify the problem by showing that only the different part needs to be considered (see Sec. 5.4), which significantly streamlines the computation of the value-to-weight ratio (see Theorem 5.6). Finally, we show that the uniform kernel reduces to the absorbing kernel when $|\mathcal{V}|\to\infty$, i.e., Figure 2(c) gradually becomes Figure 2(a), giving more theoretical insights.
Certified robustness using synonyms substitution. Ye et al. (2020) perturbs words into synonyms (including the original word) with the same probability to achieve certified robustness against word substitution attacks using synonyms. This certified bound closely resembles our DiffTextPure-Absorb method. Specifically, for any perturbed sentence $\bm{z}$, either it cannot result from perturbing the natural or adversarial sentence (trading rate of 0), or it is derived from both with the same probability (trading rate of 1). Consequently, the procedure of certifying using this synonym distribution is the same as that of our absorbing kernel. This approach cannot be generalized to certify word substitution attacks beyond synonyms, as perturbing uniformly into each word in the whole vocabulary with the same probability would completely disrupt the semantics.
Certified robustness for large language models. Kumar et al. (2023) first certify large language models against suffix attacks and insertion attacks by randomly deleting tokens. In our notation, they set $p(\bm{z}|\bm{x})$ as a uniform distribution over sentences that have deleted fewer than $k$ tokens from $\bm{x}$, and they set the threshold to infinitesimally small, i.e., as long as there is one harmful $\bm{z}$, they classify $\bm{x}$ as harmful. Therefore, their certified accuracy is exactly the empirical accuracy of detectors on the original text. However, this heuristic method has a very large false positive rate, because a benign sentence can also be deleted into a harmful one. For example, “How to prevent making a bomb” can be deleted into “How to make a bomb.” From our framework, we can easily propose a fix for their method by setting the threshold $\tau$ to a value greater than zero.
Ji et al. (2024) also use a language model as a purifier and certify large language models on classification tasks, but not on safety tasks. The most significant difference between safety and classification is that safety requires a very low false positive rate, while classification does not. Certifying safety, therefore, is much more difficult than certifying classification.
Robey et al. (2023) propose smoothing a language model by randomly perturbing each character, rather than tokens. They also do not certify their defense. Their theorem is based on an assumption they define themselves, called k-stable, which states that perturbing $k+1$ characters would result in a change. This assumption is impractical and implicitly implies robustness. In this work, we do not make any such assumptions. Instead, we certify each input $\bm{x}$ independently, rather than relying on a distribution.
B.5 On Discrete Diffusion Models
Discrete diffusion models extend traditional diffusion models to the discrete domain, enabling the modeling of language inputs (Meng et al., 2022; Campbell et al., 2022; Lou et al., 2023). Given a vocabulary $\mathcal{V}=\{1,\cdots,|V|\}$, sequence length $N$, a data distribution $p:=p_{0}\in\mathcal{V}^{N}$, the forward process creates a sequence of distributions $p_{t}$ by randomly perturbing each word according to a continuous-time Markov chain:
$\frac{dp_{t}}{dt}=Q_{t}p_{t}.$ (3)
Typically, $Q_{t}$ is defined as $\sigma(t)Q$ for simplicity, where $\sigma(t)$ is a monotonic noise schedule designed to ensure that $p_{T}$ approaches a simple prior distribution $p_{prior}$. Eq. 4 provides two frequency choices of $Q$. when $Q=Q^{\rm uniform}$, this Markov chain progressively and uniformly perturbs each word to any other word over time. When $Q=Q^{\rm absorb}$, it gradually perturbs each word into an absorbing token.
$Q^{\rm uniform}=\begin{bmatrix}1-N&1&\cdots&1\\ 1&1-N&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1-N\end{bmatrix},\quad Q^{\rm absorb}=\begin{bmatrix}-1&0&\cdots&0&0\\ 0&-1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-1&0\\ 1&1&\cdots&1&0\end{bmatrix}.$ (4)
The forward process has an analytical form due to its simplicity. For the $i$-th word $\bm{x}_{0}^{i}$, $p_{t|0}(\cdot|\bm{x}_{0}^{i})=\exp(\int_{0}^{t}\sigma(s)dsQ)_{\bm{x}_{0}^{i}}$. It also has a well-known reversal given by another diffusion matrix $\overline{Q}_{t}$ (Kelly, 2011). For the $i$-th word, the reversal is:
[…]
where $\bm{y}$ is another sentence that differs from $\bm{x}_{t}$ only at $i$-th position, $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ is referred as the concrete score. Once we train a score network $s_{\theta}(\bm{x}_{t},t)$ to approximate the concrete score, we can sample new instances using Eq. 5 by substituting the unknown score $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ with the neural network-estimated score $s_{\theta}(\bm{x},t)$ (Meng et al., 2022; Lou et al., 2023). Unlike the forward process, the reverse process lacks an analytical form due to the involvement of a neural network. Consequently, numerical methods such as an Euler solver or a $\tau$-leaping solver are typically employed to approximate the backward Markov chain.
Figure 1: Comparison between DiffPure (Nie et al., 2022) and DiffTextPure using diffusion language models.
B.6 Certification on Different Tasks
Case 1: Image Classification. In image classification, $f$ can be a classifier mapping from the image domain to one interested class in $K-1$ probability simplex. The smoothing distribution $p(\bm{z}|\bm{x})$ can be a Gaussian distribution (Cohen et al., 2019; Chen et al., 2024b), a Uniform distribution (Levine & Feizi, 2021; Lee et al., 2019), Laplacian distribution (Teng et al., 2020), or other types of distributions. If we can certify that $p_{adv}\geq 0.5$ in Definition 5.1, it guarantees that the classifier will consistently produce the correct result for all $\bm{x}_{adv}$ satisfying $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This is because the probability of the true class remains the highest among all output probabilities.
Case 2: Text Classification. Similarly, for a text classifier $f:\mathcal{V}^{N}\to[0,1]$, that maps from a text to a probability of outputting a target class that we are interested in, the smoothing distribution can be derived from the noisy process of diffusion models, $p_{t|0}(\bm{x}_{\tau}|\bm{x})$, such as randomly replacing or masking words (Lou et al., 2023), as described in Sec. 4. If we can certify that $p_{adv}\geq 0.5$ for the correct class $y$, it ensures that $y$ remains the largest output of $g(\bm{x}_{adv})$, guaranteeing robust classification.
Case 3: Text Safety. This has already been extensively discussed in Sec. 5.1.
[…]
Appendix C Proofs
C.1 Proof of Theorem 5.3
The optimality of the greedy algorithm in Theorem 5.3 has been extensively proven (Aho & Hopcroft, 1974; Cormen et al., 2022). The proof is typically conducted by contradiction. By sorting the items by their value-to-weight ratio, assume that there exists a better selection than the one obtained by selecting items based on their value-to-weight ratios. Comparing the differing items in these two selections, both must have the same volume, but the selection based on value-to-weight ratio will always have a higher ratio, and thus a higher value. Therefore, in the fractional knapsack problem, it is impossible to find a better approach than selecting items in descending order of their value-to-weight ratio.
Another proof, more closely related to the approach in Cohen et al. (2019), uses the method of Lagrange multipliers. Our goal is to find the minimal solution to a constrained optimization problem:
$\min_{f,\bm{x}_{adv}}g(\bm{x}_{adv})=\min_{f,\bm{x}_{adv}}\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})\quad\text{s.t.}\quad g(\bm{x})=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=p_{A},\,\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d.$
We construct the Lagrangian:
$\mathcal{L}=\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})+\lambda\left(\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})-p_{A}\right).$
The hypothesis set of the base function $f$ consists of all bounded functions. Normalizing them to $[0,1]$, we define the hypothesis set as =$\mathcal{F}=\{f:\mathcal{X}\to[0,1]\}$. Thus, each $f(\bm{z})$ can take any value in $[0,1]$. We treat $f(\bm{z})$ for each $\bm{z}$ as a variable and compute the derivative of $\mathcal{L}$ with respect to $f(\bm{z})$. For each $\bm{z}$, we have (total of $|\mathcal{X}|$):
$\frac{\partial\mathcal{L}}{\partial f(\bm{z})}=p(\bm{z}|\bm{x}_{adv})+\lambda p(\bm{z}|\bm{x}).$
Taking the derivative with respect to $\lambda$, we have the $|\mathcal{X}|+1$ equality:
$\sum_{\bm{z}}f(\bm{z})p(\bm{z}|\bm{x})=p_{A}.$
Since we have total $|\mathcal{X}|+1$ variables, including $|\mathcal{X}|$ for $f(\bm{z})$ and one for $\lambda$, we can solve this problem.
If $p(\bm{z}|\bm{x}_{adv})+\lambda p(\bm{z}|\bm{x})\leq 0$, i.e., $\lambda\leq-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, then $\mathcal{L}$ is a monotonically decreasing function of $f(\bm{z})$. Therefore, $f(\bm{z})$ should be set to 1. Conversely, if $p(\bm{z}|\bm{x}_{adv})+\lambda p(\bm{z}|\bm{x})\geq 0$, i.e., $\lambda\geq-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$, then $\mathcal{L}$ is a monotonically increasing function of $f(\bm{z})$. Therefore, $f(\bm{z})$ should be set to 0.
In other words, if the value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is less than $\lambda$, then $f(\bm{z})$ should be set to 0. If the value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is greater than $\lambda$, then $f(\bm{z})$ should be set to 1. Therefore, the algorithm to solve this problem is to first sort the value-to-weight ratios and then set the corresponding function values to 1 in order, until the constraint $g(\bm{x})=p_{A}$ is satisfied (which controls $\lambda$).
Remark C.1.
Further narrowing of the hypothesis set can yield better solutions for this constrained optimization problem, e.g., restricting to binary functions =$\mathcal{F}=\{f:\mathcal{X}\to\{0,1\}\}$ or functions with Lipschitz continuity (Chen et al., 2024a; Delattre et al., 2024).
C.2 Proof of Theorem 5.6
Let $v(i,j)$ be the probability measure on $p(\bm{z}|\bm{x})$ for $\{\bm{z}|p(\bm{z}|\bm{x})=\alpha^{i}\bar{\beta}^{d-i}\land p(\bm{z}|\bm{x}_{adv})=\alpha^{j}\bar{\beta}^{d-j}\}$. To calculate $v(i,j)$, we need to compute the number of items in this set and multiply by $\alpha^{i}\bar{\beta}^{d-i}$.
Since there is a $d$-token difference between $\bm{x}$ and $\bm{x}_{adv}$, $\bm{z}$ can only be derived from both $\bm{x}$ and $\bm{x}_{adv}$ if $i+j\geq d$. There are three types of tokens in $\bm{z}$:
  • Tokens that differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$.
  • Tokens that differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$.
  • Tokens that differ from both.
These tokens can appear anywhere in the adversarial part.
The first way to express this combination number is by first considering the tokens that differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. These tokens account for $\binom{d}{d-i}$. Among the remaining $i$ tokens, $i+j-d$ tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, so they contribute $\binom{i}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. Therefore, we have:
$\binom{d}{d-i}\binom{i}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}=\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}.$
Similarly, we can express this combination number from the perspective of $\bm{x}_{adv}$ instead of $\bm{x}$. First, we consider the tokens that differ from the corresponding part of $\bm{x}$ but match $\bm{x}_{adv}$. These tokens contribute $\binom{d}{d-j}$. Among the remaining $j$ tokens, $i+j-d$ tokens must differ from both $\bm{x}_{adv}$ and $\bm{x}$, contributing $\binom{j}{i+j-d}$. The remaining tokens differ from the corresponding part of $\bm{x}_{adv}$ but match $\bm{x}$. Thus, we get:
$\binom{d}{d-j}\binom{j}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}=\binom{d}{j}\binom{j}{d-i}(|\mathcal{V}|-2)^{i+j-d}.$
These two combinations are actually the same, as shown by the symmetrization lemma in Theorem C.3. This symmetry provides many favorable properties for the uniform kernel.
Below, we present three case studies to directly illustrate this combination number.
C.2.1 Case Study: $d=1$
When $d=1$, there are four types of cases:
$\bar{\beta}\to\alpha$. We use $\bar{\beta}\to\alpha$ as a more intuitive way to express the transition from $\bar{\beta}$ in $p_{A}$ to $\alpha$ in $p_{adv}$. There is only one $\bm{z}$ that satisfies this transition, which corresponds to not changing any tokens from $\bm{x}$.
$\bar{\beta}\to\bar{\beta}$. $\bm{z}$ must be same as both $\bm{x}$ and $\bm{x}_{adv}$. This is impossible.
$\alpha\to\alpha$. This means the adversarial part of $\bm{z}$ differs from both $\bm{x}$ and $\bm{x}_{adv}$. There are $|\mathcal{V}|-2$ possible $\bm{z}$ that satisfy this condition.
$\alpha\to\bar{\beta}$. There is only one $\bm{z}$ that satisfies this condition, and it must be identical to $\bm{x}_{adv}$.
C.2.2 Case Study: $d=2$
When $d=2$, there are $93^{2}=$ 9 cases.
$\bar{\beta}^{2}\to\alpha^{2}$. There is only one $\bm{z}$ that satisfies this condition, and it must be identical to $\bm{x}$.
$\bar{\beta}^{2}\to\bar{\beta}\alpha$. This is the case where $\bm{z}$ is the same as $\bm{x}$, but differs from $\bm{x}_{adv}$ by only one token. This case is impossible.
$\bar{\beta}^{2}\to\bar{\beta}^{2}$. This is the case where $\bm{z}$ is the same as $\bm{x}$, but differs from $\bm{x}_{adv}$ by two tokens. This case is also impossible.
$\bar{\beta}\alpha\to\alpha^{2}$. One token must be the same as $\bm{x}$, while the other must differ from both $\bm{x}$ and $\bm{x}_{adv}$. There are $\binom{2}{1}(|\mathcal{V}|-2)$ possible $\bm{z}$ that satisfy this condition.
$\bar{\beta}\alpha\to\bar{\beta}\alpha$. One token must be the same as $\bm{x}$, while the other must be the same as $\bm{x}_{adv}$. There are $\binom{2}{1}=2$ possible $\bm{z}$ that satisfy this condition.
$\bar{\beta}\alpha\to\bar{\beta}^{2}$. This is the case where $\bm{z}$ is the same as $\bm{x}_{adv}$, but differs from $\bm{x}$ by one token. This case is impossible.
$\alpha^{2}\to\alpha^{2}$. All tokens must differ from both $\bm{x}$ and $\bm{x}_{adv}$. There are $(|\mathcal{V}|-2)^{2}$ possible $\bm{z}$ that satisfy this condition.
$\alpha^{2}\to\bar{\beta}\alpha$. One token must be the same as $\bm{x}_{adv}$, and the other must differ from both $\bm{x}$ and $\bm{x}_{adv}$. There are $\binom{2}{1}(|\mathcal{V}|-2)$ possible $\bm{z}$ that satisfy this condition.
$\alpha^{2}\to\bar{\beta}^{2}$. This case requires $\bm{z}$ to be identical to $\bm{x}_{adv}$. There is only one such $\bm{z}$.
From this case study, we can see that although there are $(d+1)^{2}$ cases since both $i$ and $j$ have $d+1$ choices, we only need to consider $i+j\geq d$. If $i+j<d$, then no $\bm{z}$ can satisfy this condition.
C.2.3 Case study: $d=3$
We enumerate all cases following the previous order.
$\bar{\beta}^{3}\to\alpha^{3}$. There is only one $\bm{z}$ that satisfies this condition, and it must be identical to $\bm{x}$.
$\bar{\beta}^{2}\alpha\to\alpha^{3}$. Two tokens must be the same as $\bm{x}$, and one token should differ from both. There are $\binom{3}{1}(|\mathcal{V}|-2)$ $\bm{z}$.
$\bar{\beta}^{2}\alpha\to\bar{\beta}\alpha^{2}$. Two tokens must be the same as $\bm{x}$, and one token must be the same as $\bm{x}_{adv}$. There are $\binom{3}{1}=3$ $\bm{z}$.
$\bar{\beta}\alpha^{2}\to\alpha^{3}$. One token must be the same as $\bm{x}$, and the other two tokens should differ from both. There are $\binom{3}{1}(|\mathcal{V}|-2)^{2}$ $\bm{z}$.
$\bar{\beta}\alpha^{2}\to\bar{\beta}\alpha^{2}$. One token must be the same as $\bm{x}$, one token must be the same as $\bm{x}_{adv}$, and one token should differ from both. There are $\binom{3}{1}\binom{2}{1}(|\mathcal{V}|-2)$ $\bm{z}$.
$\bar{\beta}\alpha^{2}\to\bar{\beta}^{2}\alpha$. One token must be the same as $\bm{x}$, and two tokens must be the same as $\bm{x}_{adv}$. There are $\binom{3}{1}=3$ $\bm{z}$.
$\alpha^{3}\to\alpha^{3}$. All tokens should differ from both. There are $(|\mathcal{V}|-2)^{3}$ $\bm{z}$.
$\alpha^{3}\to\bar{\beta}\alpha^{2}$. One token must be the same as $\bm{x}_{adv}$, and two tokens should differ from both. There are $\binom{3}{1}(|\mathcal{V}|-2)^{2}$ $\bm{z}$.
$\alpha^{3}\to\bar{\beta}^{2}\alpha$. Two tokens must be the same as $\bm{x}_{adv}$, and one token should differ from both. There are $\binom{3}{2}(|\mathcal{V}|-2)$ $\bm{z}$.
$\alpha^{3}\to\bar{\beta}^{3}$. The result must be identical to $\bm{x}_{adv}$. Only one $\bm{z}$.
C.3 Proof of Theorem 5.5
The volume of $L_{1}$ can be simplified as follows:
$\sum_{\bm{z}\in L_{1}}p(\bm{z}|\bm{x})=\sum_{i=d}^{N}\binom{N}{i}\beta^{i}\bar{\beta}^{N-i}\frac{\binom{N-d}{i-d}}{\binom{N}{i}}=\sum_{i=d}^{N}\binom{N-d}{i-d}\beta^{i}\bar{\beta}^{N-i}$
$=$ $\sum_{i=0}^{N-d}\binom{N-d}{i}\beta^{i+d}\bar{\beta}^{N-d-i}=\beta^{d}\sum_{i=0}^{N-d}\binom{N-d}{i}\beta^{i}\bar{\beta}^{N-d-i}=\beta^{d}.$
Accordingly, the volume of $L_{2}$ is:
$\sum_{\bm{z}\in L_{2}}p(\bm{z}|\bm{x})=1-\sum_{\bm{z}\in L_{1}}p(\bm{z}|\bm{x})=1-\beta^{d}.$
This simple result enables us to intuitively illustrate the greedy algorithm using $p_{adv}-p_{A}$ graph. See Sec. C.4 and Figure 2(a) for detail.
One can also interpret the certified bound for absorbing kernel in another way, similar to Zeng et al. (2023):
For absorbing kernel, the region of smoothed examples $\bm{z}\sim p(\cdot|\bm{x})$ can be divided into two parts. The first part, $L_{1}$, consists of cases where the forward process has masked all adversarial tokens. These samples can also be generated from $p(\cdot|\bm{x}_{adv})$.
The second part, $L_{2}$, includes cases where none of the adversarial tokens are masked. The smoothed input $\bm{z}$ in this case cannot be derived from either $p(\cdot|\bm{x}_{adv})$ or $p(\cdot|\bm{x})$.
In the worst-case scenario for adversarial input, all tokens in the adversarial suffix differ from those in the original input. If any token in the suffix of $\bm{x}$ matches that of $\bm{x}_{adv}$, then it cannot be obtained from $p(\cdot|\bm{x}_{adv})$, and vice versa. Clearly, $L_{1}\cup L_{2}=\mathcal{V}^{N}$.
Therefore, the output $g(\bm{x}_{adv})$ must satisfy $g(\bm{x}_{adv})\geq\sum_{\bm{z}\in L_{1}}f(\bm{z})p(\bm{z}|\bm{x}_{adv})$. Note that for $\bm{z}\in L_{1}$, $p(\bm{z}|\bm{x}_{adv})=p(\bm{z}|\bm{x})$, so Theorem 5.5 holds. Additionally, there exists a worst-case $f$ where $f=0$ for all $\bm{z}\in L_{2}$, making this bound tight.
C.4 Analytic Solution of Certified Robustness using Absorbing Kernel
(a) $p_{adv}-p_{A}$ plot for the absorbing kernel.
(b) $p_{adv}-p_{A}$ plot for the uniform kernel.
(c) $p_{adv}-p_{A}$ plot for different vocabulary size.
Figure 2: Comparison of $p_{adv}-p_{A}$ plots for the absorbing kernel and the uniform kernel, illustrating the Knapsack algorithm. $p_{adv}$ is plotted on the vertical axis, and $p_{A}$ on the horizontal axis. When the vocabulary size $|\mathcal{V}|$ increases, the $p_{adv}-p_{A}$ of the uniform kernel gradually shifts downward and to the right, eventually matching that of the absorbing kernel.
We analyze the $p_{adv}-p_{A}$ plots (where $p_{adv}$ is on the vertical axis and $p_{A}$ is on the horizontal axis), which provide a direct illustration of the Knapsack algorithm. As shown in Figure 2(a), $p_{adv}=0$ when $p_{A}\leq 1-\beta^{d}$. When $p_{A}\geq 1-\beta^{d}$, we trade $p_{adv}$ for $p_{A}$ at a trading rate of 1 (indicated by a slope of 1).
To achieve certification, $p_{adv}$ must exceed $\tau$. This requires $p_{A}\geq 1-\beta^{d}+\tau$. Solving for $d$, we derive:
$p_{A}\geq 1-\beta^{d}+\tau\Leftrightarrow\beta^{d}\geq 1-p_{A}+\tau\Leftrightarrow d\log\beta\geq\log(1-p_{A}+\tau)\Leftrightarrow d\leq\frac{\log(1-p_{A}+\tau)}{\log\beta}.$
This means the certified radius of absorbing kernel is $\lfloor\frac{\log(1-p_{A}+\tau)}{\log\beta}\rfloor$.
We do not use this analytic solution in this paper, since running the knapsack solver and using this analytic solution both require $O(1)$ time complexity.
C.5 Proof of Theorem 5.7
Since the certified robustness of the uniform kernel does not have an analytic solution, proving Theorem 5.7 requires some subtle observations.
Notice that for the absorbing kernel, $p_{adv}=g(\bm{x}_{adv})=0$ when $p_{A}\leq 1-\beta^{d}$, and it increases linearly with $p_{A}$ with a slope of 1, as the value-to-weight ratio is 1 (when all $\bm{z}_{s}$ are mask tokens, $p(\bm{z}|\bm{x})=p(\bm{z}|\bm{x}_{adv})=\beta^{d}$). Therefore, when trading $p_{adv}$ with $p_{A}$, the trading rate (value-to-weight ratio) is either 0 or 1, with 0 occurring first and 1 following.
Think about the $p_{adv}-p_{A}$ plots (where $p_{adv}$ is on the vertical axis and $p_{A}$ is on the horizontal axis). If we can prove that once we begin using a trading rate of 1 in the absorbing kernel, we are already using a trading rate greater than 1 in the uniform kernel, we can conclude that the $p_{adv}$ for the uniform kernel will always be greater than that for the absorbing kernel. Consequently, when using the same threshold $\tau$, the certified radius for the uniform kernel will always outperform that of the absorbing kernel.
Formally, we want to prove that:
$\sum_{i<j,i+j\geq d}v(i,j)\leq 1-\beta^{d}.$ (6)
The right-hand side represents the starting point for the absorbing kernel when using a trading rate of $1$, and the left-hand side represents the starting point for the uniform kernel with the same trading rate. This is because, when $i<j$, the value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is given by
$\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}\leq 1.$
The condition $\left(\frac{\alpha}{\bar{\beta}}\right)\leq 1$ is equivalent to $\bar{\beta}\geq\frac{1}{\mathcal{V}}$, and this is always satisfied because at $t_{\max}$ the uniform prior assigns equal probability $\frac{1}{\mathcal{V}}$ to all tokens. Therefore, Eq. 6 provides a sufficient condition for Theorem 5.7.
In the following subsections, we first present a complete proof of Eq. 6. Then, we analyze some simple cases to provide intuition on how we arrive at this proof.
C.5.1 Final Proof of Sufficient Condition Eq. 6
We first give the following lemma:
Lemma C.2.
The summation of $v(i,j)$ over all valid $i,j$ equals $1$, i.e.,
$\sum_{i+j\geq d}v(i,j)=1.$
Proof.
The above lemma is expected since $v(i,j)$ represents a probability measure over $i,j$. We prove this by the following transformations:
$\sum_{i+j\geq d}\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}\bar{\beta}^{d-i}=\sum_{i=0}^{d}\sum_{j=d-i}^{d}\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}\bar{\beta}^{d-i}$
$=$ $\sum_{i=0}^{d}\sum_{j=0}^{i}\binom{d}{i}\binom{i}{j}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-2)^{i-j}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-2)^{i}\sum_{j=0}^{i}\binom{i}{j}(|\mathcal{V}|-2)^{-j}$
$=$ $\sum_{i=0}^{d}\binom{d}{i}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-2)^{i}(1+\frac{1}{|\mathcal{V}|-2})^{i}=\sum_{i=0}^{d}\binom{d}{i}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-2)^{i}(\frac{|\mathcal{V}|-1}{|\mathcal{V}|-2})^{i}$
$=$ $\sum_{i=0}^{d}\binom{d}{i}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-1)^{i}=\sum_{i=0}^{d}\binom{d}{i}\bar{\beta}^{d-i}[\alpha(|\mathcal{V}|-1)]^{i}$
$=$ $\sum_{i=0}^{d}\binom{d}{i}\bar{\beta}^{d-i}\beta^{i}=(\bar{\beta}+\beta)^{d}=1.$
Using this lemma, we upper bound Eq. 6 by:
$\sum_{i<j,i+j\geq d}v(i,j)$ $=\sum_{i+j\geq d}v(i,j)-\sum_{i\geq j,i+j\geq d}v(i,j)<1-\sum_{i=d}^{d}\sum_{j=0}^{d}v(i,j)$
$=1-\sum_{j=0}^{d}\binom{d}{d-j}(|\mathcal{V}|-2)^{j}\alpha^{d}=1-\sum_{j=0}^{d}\binom{d}{j}(|\mathcal{V}|-2)^{j}\alpha^{d}$
$=1-(|\mathcal{V}|-1)^{d}\alpha^{d}=1-[\alpha(|\mathcal{V}|-1)]^{d}=1-\beta^{d}.$
Which completes the proof of Eq. 6. Since Eq. 6 is a sufficient condition of Theorem 5.7, this also completes the proof of Theorem 5.7.
The above inequality is nearly tight. As $|\mathcal{V}|\to\infty$, the inequality approaches equality. Refer to the case study in the next section for further details.
C.5.2 Simple case study: $|\mathcal{V}|\to\infty$
In this subsection, we show that when the vocabulary size $|\mathcal{V}|\to\infty$, the above inequality approaches equality. In other words,
$\lim_{|\mathcal{V}|\to\infty}\sum_{i<j,i+j\geq d}v(i,j)=\lim_{|\mathcal{V}|\to\infty}\sum_{i<j,i+j\geq d}\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}\bar{\beta}^{d-i}=1-\beta^{d}.$
The key insight here is that $\alpha^{i}=\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}$, contain a high order term $\frac{1}{(|\mathcal{V}|-1)^{i}}$. We know that $i+j-d\leq i$ since $j\leq d$. When $i+j-d<i$, $(|\mathcal{V}|-2)^{i+j-d}\alpha^{i}=(|\mathcal{V}|-2)^{i+j-d}\frac{\beta^{i}}{(|\mathcal{V}|-1)^{i}}\to 0$. Hence, we only need to consider $i+j-d=i$, or equivalently, $j=d$. Therefore, we have the following:
$\lim_{|\mathcal{V}|\to\infty}\sum_{i<j,i+j\geq d}v(i,j)=\lim_{|\mathcal{V}|\to\infty}\sum_{i<d,i\geq 0}v(i,d)=1-\lim_{|\mathcal{V}|\to\infty}v(d,d)$
$=$ $1-\lim_{|\mathcal{V}|\to\infty}(|\mathcal{V}|-2)^{d}\alpha^{d}=1-\lim_{|\mathcal{V}|\to\infty}(|\mathcal{V}|-2)^{d}\frac{\beta^{d}}{(|\mathcal{V}|-1)^{d}}=1-\beta^{d}.$
Intuitively, the certified robustness would be the smallest when $|\mathcal{V}|\to\infty$. This inspired us to bound Eq. 6 using $j=d$. However, the last step $\frac{(|\mathcal{V}|-2)^{d}}{(|\mathcal{V}|-1)^{d}}=1$ does not hold when $|\mathcal{V}|\neq\infty$. Therefore, we consider loosing by $i=d$ when proving Eq. 6. This case study also demonstrates that Eq. 6 is almost tight since it becomes equality when $|\mathcal{V}|\to\infty$.
When $|\mathcal{V}|\to\infty$, the value-to-weight ratio $\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=\frac{\alpha^{i-j}}{\bar{\beta}^{i-j}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{i-j}$ only have three possible values: 0 when $i>j$, 1 when $i=j$, $\infty$ when $i<j$. Since for all $p_{A}\leq 1-\beta^{d}$, we have $i>j$, thus $p_{adv}=0$ for all $p_{A}\leq 1-\beta^{d}$. By symmetrization lemma (Theorem C.3), $i=j$ must hold for all $p_{A}\geq 1-\beta^{d}$. Therefore, the $p_{adv}-p_{A}$ graph of the uniform kernel and absorbing kernel is exactly the same. This means Figure 2(b) gradually goes to Figure 2(a) when $|\mathcal{V}|\to\infty$.
C.5.3 Simple case study: $d$=1,2,3
When $d=1$, the summation of volume for trading rate less than one is exactly $1-\beta$:
$\sum_{0\leq i<j\leq d}v(i,j)=v(0,1)=\bar{\beta}=1-\beta.$
When $d=2$, we have:
$\sum_{0\leq i<j\leq d}v(i,j)=v(0,1)+v(0,2)+v(1,2)=v(0,2)+v(1,2)=\bar{\beta}^{2}+2(|\mathcal{V}|-2)\bar{\beta}\alpha$
$=$ $(1-\beta)^{2}+2(1-\beta)\beta\frac{|\mathcal{V}|-2}{|\mathcal{V}|-1}\leq 1-2\beta+\beta^{2}+2(1-\beta)\beta=1-\beta^{2}.$
When $d=3$, the inequality $(|\mathcal{V}|-2)\alpha\leq\beta$ becomes too loose. Thus, we need to prove this in a slightly more refined way:
$\sum_{0\leq i<j\leq d}v(i,j)=v(0,3)+v(1,3)+v(1,2)+v(2,3)=\bar{\beta}^{3}+3(|\mathcal{V}|-2)\bar{\beta}^{2}\alpha+3\bar{\beta}^{2}\alpha+3(|\mathcal{V}|-2)^{2}\bar{\beta}\alpha^{2}$
$=$ $\bar{\beta}^{3}+3(|\mathcal{V}|-1)\bar{\beta}^{2}\alpha+3(|\mathcal{V}|-2)^{2}\bar{\beta}\alpha^{2}\leq(1-\beta)^{3}+3(1-\beta)^{2}\beta+3(1-\beta)\beta^{2}$
$=$ $(1-\beta)^{3}+3(1-\beta)\beta=1-3\beta+3\beta^{2}-\beta^{3}+3\beta-3\beta^{2}=1-\beta^{3}.$
This motivate us to provide the general proof in Eq. 6
C.6 Symmetrization Lemma
In this section, we present the symmetrization lemma for the uniform kernel. This lemma provides an intuitive understanding of the $p_{adv}-p_{A}$ graph for the uniform kernel and plays a crucial role in several theorems presented in this paper.
Theorem C.3.
The $p_{adv}-p_{A}$ graph of the uniform kernel is symmetric with respect to the line $p_{adv}=-p_{A}+1$.
Proof.
We prove this theorem in three steps.
Symmetrization of the slope:
The $p_{adv}-p_{A}$ graph is a piecewise linear function. We begin by proving that if there exists a linear segment with slope $k$, there must also be a corresponding linear segment with slope $\frac{1}{k}$.
This is evident because the trading rate, given by
$\frac{\alpha^{j}\bar{\beta}^{d-j}}{\alpha^{i}\bar{\beta}^{d-i}}=\left(\frac{\alpha}{\bar{\beta}}\right)^{j-i},$
can only take $12d+$ 1 distinct values, specifically:
$\left\{\left(\frac{\alpha}{\bar{\beta}}\right)^{-d},\dots,\left(\frac{\alpha}{\bar{\beta}}\right)^{-1},1,\left(\frac{\alpha}{\bar{\beta}}\right)^{1},\dots,\left(\frac{\alpha}{\bar{\beta}}\right)^{d}\right\}.$
Thus, the slope must exhibit symmetry.
Symmetry of each line segment with respect to the x-axis and y-axis:
In other words, we need to prove that if a line segment with slope $k$ trades $B$ of $p_{adv}$ using $A$ of $p_{A}$, then the line segment with slope $\frac{1}{k}$ must trade $A$ of $p_{adv}$ using $B$ of $p_{A}$.
Consider the part of the graph where
$\{p(\bm{z}|\bm{x})=\alpha^{i}\bar{\beta}^{d-i}\land p(\bm{z}|\bm{x}_{adv})=\alpha^{j}\bar{\beta}^{d-j}\},$
which trades $v(i,j)$ of $p_{A}$ for
$v(i,j)\cdot\frac{\alpha^{j}\bar{\beta}^{d-j}}{\alpha^{i}\bar{\beta}^{d-i}}\text{ of }p_{adv}.$
For the symmetric case,
$\{p(\bm{z}|\bm{x})=\alpha^{j}\bar{\beta}^{d-j}\land p(\bm{z}|\bm{x}_{adv})=\alpha^{i}\bar{\beta}^{d-i}\},$
the trade is $v(j,i)$ of $p_{A}$ for
$v(j,i)\cdot\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}\text{ of }p_{adv}.$
Thus, we only need to prove the following two equalities:
$v(i,j)\cdot\frac{\alpha^{j}\bar{\beta}^{d-j}}{\alpha^{i}\bar{\beta}^{d-i}}=v(j,i),$
and
$v(j,i)\cdot\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=v(i,j).$
We prove the first equality as follows:
$\quad v(i,j)\cdot\frac{\alpha^{j}\bar{\beta}^{d-j}}{\alpha^{i}\bar{\beta}^{d-i}}$
$=\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{i}\bar{\beta}^{d-i}\cdot\frac{\alpha^{j}\bar{\beta}^{d-j}}{\alpha^{i}\bar{\beta}^{d-i}}$
$=\binom{d}{i}\binom{i}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$
$=\binom{d}{i}\binom{i}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-j}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}\binom{B}{C}=\binom{A}{C}\binom{A-C}{B-C}$
$=\binom{d}{i+j-d}\binom{2d-i-j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=\binom{d}{j}\binom{j}{i+j-d}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{C}\binom{A-C}{B-C}=\binom{A}{B}\binom{B}{C}$
$=\binom{d}{j}\binom{j}{d-i}(|\mathcal{V}|-2)^{i+j-d}\cdot\alpha^{j}\bar{\beta}^{d-j}$ by ($\text{by }\binom{A}{B}=\binom{A}{A-B}$
$=v(j,i)$
The second equality can be proven in a similar manner. Alternatively, one can simply swap all occurrences of $i$ and $j$ in the first equality, which directly yields the second equality. Specifically, by replacing $i\leftrightarrow j$, we get the following:
$v(j,i)\cdot\frac{\alpha^{i}\bar{\beta}^{d-i}}{\alpha^{j}\bar{\beta}^{d-j}}=v(i,j),$
which is the second equality that we aimed to prove.
Symmetry of Endpoints of Each Segment:
From left to right, the trading rate (slope) increases, while from right to left, the slope decreases.
The first point $(0,0)$ corresponds to $(1,1)$. Then, the minimal slope trade occurs when $A_{1}$ of $p_{A}$ is traded for $B_{1}$ of $p_{adv}$, and the maximum slope trade occurs when $B_{1}$ of $p_{A}$ is traded for $A_{1}$ of $p_{adv}$. Thus, the points $(A_{1},B_{1})$ and $(1-B_{1},1-A_{1})$ lie on the graph.
Now, assume that the first $m$ points are symmetric. Thus, the points $\left(\sum_{i=1}^{m}A_{i},\sum_{i=1}^{m}B_{i}\right)$ and $\left(1-\sum_{i=1}^{m}B_{i},1-\sum_{i=1}^{m}A_{i}\right)$ are on the graph.
On the left $(m+1)$-th segment, we trade $A_{m+1}$ of $p_{A}$ for $B_{m+1}$ of $p_{adv}$, and on the right side, we trade $B_{m+1}$ of $p_{A}$ for $A_{m+1}$ of $p_{adv}$. Thus, the points $\left(\sum_{i=1}^{m+1}A_{i},\sum_{i=1}^{m+1}B_{i}\right)$ and $\left(1-\sum_{i=1}^{m+1}B_{i},1-\sum_{i=1}^{m+1}A_{i}\right)$ are also on the graph.
By induction, this symmetry holds for all subsequent segments. Therefore, all endpoints of this piecewise linear function are symmetric, and hence, the entire $p_{adv}-p_{A}$ graph is symmetric.
An illustration of $p_{adv}-p_{A}$ graph using uniform kernel is presented in Figure 2(b).
Through the symmetrization lemma Theorem C.3, we have the following corollary, which will be used in Sec. C.7.
Corollary C.4.
The $p_{adv}-p_{A}$ plot intersects the axis of symmetry $p_{adv}=-p_{A}+1$ at the part with slope 1.
Proof.
This can be easily proved by contradiction. If the intersection part has a slope other than 1, let us assume it is $k$. Then, the slope of 1 must be either to the left or right of the axis of symmetry. Due to the symmetry, the other side must still have a slope of 1. Since the slope is a non-decreasing function of $p_{A}$, this implies that $11<k<$ 1, which leads to a contradiction. Therefore, this corollary is true. ∎
C.7 Relationship Between $|\mathcal{V}|$ and Certified Radius
We propose the following conjecture:
Conjecture C.5.
The certified robustness of the uniform kernel is a decreasing function of $|\mathcal{V}|$. Formally, given the same accuracy $p_{A}$, threshold $\tau$, and perturbing probability $\beta$, for $|\mathcal{V}_{1}|\geq|\mathcal{V}_{2}|$, we have:
$\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V}_{1})\leq\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V}_{2}).$
This conjecture is reasonable because, as the vocabulary size increases, the input space also increases. Some studies suggest that the existence of adversarial examples arises from the exponentially large input space.
However, we have not been able to prove this conjecture. Instead, we propose a weaker version of this conjecture, which can be easily proved:
Theorem C.6.
There exists a constant $C_{\mathcal{V}}$ such that, given the same accuracy $p_{A}$, threshold $\tau$, and perturbing probability $\beta$, for $|\mathcal{V}_{1}|\geq|\mathcal{V}_{2}|>C_{\mathcal{V}}$, we have:
$\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V}_{1})\leq\text{certify}(\text{uniform},p_{A},\tau,\beta,\mathcal{V}_{2}).$
In other words, the certified radius is a decreasing function when $|\mathcal{V}|\geq C_{\mathcal{V}}$. This constant can be bounded by:
$C_{\mathcal{V}}\leq d+1.$
Using the symmetrization lemma (Theorem C.3), we only need to prove the case where the trading rate $\left(\frac{\alpha}{\bar{\beta}}\right)^{j-i}\leq 1$, i.e., $j\geq i$. In this proof, unless stated otherwise, we assume $j\geq i$.
First, notice that the trading rate $\left(\frac{\alpha}{\bar{\beta}}\right)^{j-i}$ is monotonically decreasing as $|\mathcal{V}|$ increases. Following the notation from the previous section, let $A_{k}$ denote the $k$-th minimal $v(i,j)$, and let $B_{k}$ represent the trading result using $A_{k}$. The endpoints of each piecewise linear function are given by $\left(\sum_{i=1}^{m+1}A_{i},\sum_{i=1}^{m+1}B_{i}\right)$. As long as we can show that $\sum_{i=1}^{m+1}A_{i}$ is monotonically increasing as $|\mathcal{V}|$ grows for every $m$, we can apply induction to demonstrate that for every endpoint, $p_{adv}(p_{A},\mathcal{V}_{1})\leq p_{adv}(p_{A},\mathcal{V}_{2})$. This will establish that the inequality holds at every point, completing the proof.
Proof.
Step 1. $v(i,j)$ is a monotonically increasing function of $|\mathcal{V}|$ when $|\mathcal{V}|\geq d+1\geq C_{\mathcal{V}}$:
Lets assume $|\mathcal{V}_{1}|\geq|\mathcal{V}_{2}|$. Denote $A_{i}(\mathcal{V})$ as the volume of $i$-th minimal trading rate. $B_{i}(\mathcal{V})$ as the corresponding volume times the trading rate. Let $r=j-i$. For the same $r$, we have the same trading rate. We calculate $\sum_{i=1}^{m}A_{i}$ by summing $v(i,i+r)$ in the order of $r$ (i.e., from larger trading rates to smaller trading rates):
$\sum_{i=1}^{m}A_{i}=\sum_{r=d}^{r(m)}\sum_{i=0}^{d-r}v(i,i+r),$
where $r(m)$ is an integer that controls the total number of summations equal to $m$. We can rewrite this summation as:
$\sum_{i=1}^{m}A_{i}=\sum_{i=0}^{i(m)}\sum_{j=d}^{j(i,m)}v(i,j).$
From Lemma C.2, we have:
$\sum_{i=0}^{i(m)}\sum_{j=d}^{0}v(i,j)=\sum_{i=0}^{i(m)}\binom{d}{i}\alpha^{i}\bar{\beta}^{d-i}(|\mathcal{V}|-1)^{i}=\sum_{i=0}^{i(m)}\binom{d}{i}\beta^{i}\bar{\beta}^{d-i}.$
This is independent of $|\mathcal{V}|$. Since:
$\sum_{i=0}^{i(m)}\sum_{j=d}^{j(i,m)}v(i,j)=\sum_{i=0}^{i(m)}\sum_{j=d}^{0}v(i,j)-\sum_{i=0}^{i(m)}\sum_{j=0}^{j(i,m)-1}v(i,j),$
and for $j<d$, we have $i+j-d<i$, thus the volume
$v(i,j)=\binom{d}{i}\binom{i}{d-j}\cdot\beta^{i}\frac{(|\mathcal{V}|-2)^{i+j-d}}{(|\mathcal{V}|-1)^{i}}\bar{\beta}^{d-i}$
has a higher order term in the denominator than in the numerator. Therefore, there exists a constant $C_{\mathcal{V}}$ such that for all $|\mathcal{V}|\geq C_{\mathcal{V}}$, this is a monotonically decreasing function of $|\mathcal{V}|$.
Obviously, this constant can be bounded by:
$C_{\mathcal{V}}\leq\max_{C_{x},a,b}\text{ such that }\frac{(x-2)^{a}}{(x-1)^{b}}\text{ for }0\leq a<b\leq d\text{ is a monotonically decreasing function when }x>C_{x}.$
Taking the derivative with respect to $x$, setting it to zero:
$\frac{a(x-2)^{a-1}(x-1)^{b}-b(x-1)^{b-1}(x-2)^{a}}{(x-1)^{2}b}<0\;\Leftrightarrow\;x>\max_{a,b}\frac{2b-a}{b-a}=\max_{a,b}1+\frac{b}{b-a}=1+d.$
Therefore, we have:
$C_{\mathcal{V}}\leq d+1.$
A constant function of $|\mathcal{V}|$ minus a monotonically decreasing function of $|\mathcal{V}|$ results in a monotonically increasing function of $|\mathcal{V}|$. Thus, we conclude that $\sum_{i=1}^{m}A_{i}$ is a monotonically increasing function of $|\mathcal{V}|$ when $|\mathcal{V}|>C_{\mathcal{V}}$.
Step 2: Proof by Induction
For the first point $(A_{1},B_{1})$, as $|\mathcal{V}|$ increases, the slope of this part becomes smaller, and $A_{1}$ also increases. For all $0\leq p_{A}\leq A_{1}(\mathcal{V}_{2})$, we have $p_{adv}(p_{A},\mathcal{V}_{2})\geq p_{adv}(p_{A},\mathcal{V}_{1})$. For all $A_{1}(\mathcal{V}_{2})\leq p_{A}\leq A_{1}(\mathcal{V}_{1})$, since $\mathcal{V}_{2}$ has a higher slope, we also have $p_{adv}(p_{A},\mathcal{V}_{2})\geq p_{adv}(p_{A},\mathcal{V}_{1})$.
Now, let’s assume that the inequality $p_{adv}(p_{A},\mathcal{V}_{2})\geq p_{adv}(p_{A},\mathcal{V}_{1})$ holds for all $0\leq p_{A}\leq\sum_{i=1}^{k}A_{i}(\mathcal{V}_{1})$ for some $k$. We aim to prove that this still holds for $k+1$. For all $\sum_{i=1}^{k}A_{i}(\mathcal{V}_{1})\leq p_{A}\leq\sum_{i=1}^{k+1}A_{i}(\mathcal{V}_{1})$, the slope for $\mathcal{V}_{2}$ is always greater than or equal to that of $\mathcal{V}_{1}$, because $\sum_{i=1}^{k}A_{i}(\mathcal{V}_{1})\geq\sum_{i=1}^{k}A_{i}(\mathcal{V}_{2})$ and $\sum_{i=1}^{k+1}A_{i}(\mathcal{V}_{1})\geq\sum_{i=1}^{k+1}A_{i}(\mathcal{V}_{2})$. Since the starting points are also larger, it follows that the inequality $p_{adv}(p_{A},\mathcal{V}_{2})\geq p_{adv}(p_{A},\mathcal{V}_{1})$ still holds for all $0\leq p_{A}\leq\sum_{i=1}^{k+1}A_{i}(\mathcal{V}_{1})$. This completes the proof.
Figure 2(c) illustrates the proof idea. We are using induction to prove that the blue point is always on the right side of the corresponding red point when the trading rate is less than 1.
C.8 0-1 Knapsack
Algorithm 3 0-1 Knapsack Solver for Randomized Smoothing on Any Distribution (Dynamic Programming)
0:  Probability distributions $p(\bm{z}|\bm{x})$ and $p(\bm{z}|\bm{x}_{adv})$, output at clean example $p_{A}$, threshold $\tau$
0:  Whether $g$ is provably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$.
1:  Let $n$ be the number of items
2:  Initialize DP table $dp[i][w]=-\infty$ for all $i\leq n$ and $w\leq p_{A}$
3:  Set $dp[0][w]=0$ for all $w$, as no items result in zero value
4:  for each item $\bm{z}^{(i)}$ from $1$ to $n$ do
5:     for each possible weight $w\leq p_{A}$ do
6:        Update DP table:
$dp[i][w]\leftarrow\max(dp[i-1][w],dp[i-1][w-p(\bm{z}^{(i)}|\bm{x})]-p(\bm{z}^{(i)}|\bm{x}_{adv}))$
7:     end for
8:  end for
9:  Let $V_{\text{max}}=-\max(dp[n][w])$ for $w=0,1,\dots,W_{\max}$Maximum value obtainable with capacity $W_{\max}$}
10:  Return: $\mathbb{I}\{V_{\text{max}}\geq\tau\}$Return 1 if value $V_{\text{max}}$ is greater than or equal to threshold $\tau$, else return 0}
In Sec. 5.2, we mentioned the connection between the randomized smoothing problem and the 0-1 Knapsack problem. Specifically, if we restrict the hypothesis set of the function $f$ to hard functions that only output binary values (i.e., functions that map to $\{0,1\}$), then the problem at hand becomes a 0-1 Knapsack problem. This restriction leads to a more efficient solution where we can apply dynamic programming to obtain a tighter bound on the robustness of the function.
Let us now formalize the problem and provide a dynamic programming solution.
Given a probability distribution $p(\bm{z}|\bm{x})$ that represents the weight (or quality) of each item, and a corresponding adversarial distribution $-p(\bm{z}|\bm{x}_{adv})$ that represents the value (or profit) of each item, we are tasked with selecting a subset of items such that the total weight (i.e., the total probability mass at the clean example) does not exceed a given threshold $p_{A}$. The goal is to maximize the total value, which is the sum of the negative log-probabilities from the adversarial distribution.
This scenario naturally translates into the 0-1 Knapsack problem, where weights are given by $p(\bm{z}|\bm{x})$, values are given by $-p(\bm{z}|\bm{x}_{adv})$, the capacity of the knapsack is $p_{A}$, and the objective is to maximize the total value, subject to the constraint on the total weight.
To solve the 0-1 Knapsack problem efficiently, we employ dynamic programming (DP). The idea is to construct a DP table that tracks the maximum value that can be achieved for each possible total weight, up to the capacity $p_{A}$. The state transitions in the DP table depend on whether we include each item in the knapsack or not.
The dynamic programming solution is demonstrated in Algorithm 3. It first define $dp[i][w]$ to be the maximum value that can be obtained by considering the first $i$ items, with a knapsack capacity of $w$. For each item $\bm{z}_{i}$, if we can add it to the knapsack (i.e., if the current weight $w$ is greater than or equal to the weight of the item $p(\bm{z}_{i}|\bm{x})$), we update the DP table by considering both the inclusion and exclusion of the item:
$dp[i][w]=\max(dp[i-1][w],dp[i-1][w-p(\bm{z}_{i}|\bm{x})]-p(\bm{z}_{i}|\bm{x}_{adv})).$
This ensures that at each step, we are choosing the maximum value that can be achieved by either including or excluding the current item. After filling the DP table, the maximum value obtainable with the given capacity $p_{A}$ is the maximum value found in the last row of the table, i.e., $V_{\text{max}}=\max(dp[n][w])$ for all $w\in[0,p_{A}]$.
Finally, we check whether the maximum value obtained is greater than or equal to the threshold $\tau$. If $-V_{\text{max}}\geq\tau$, then we can certify that the function is provably robust for all distributions with $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. Otherwise, the function does not meet the robustness criterion.
The time complexity of the dynamic programming algorithm is $O(n\times n_{p_{A}})$, where $n$ is the number of items and $n_{p_{A}}$ is the number of weights that selected items can take. This is a typical time complexity for solving the 0-1 Knapsack problem using dynamic programming.
Appendix D Tricks
D.1 Parameterizing $t$ as $1-\bar{\beta}$
In the diffusion process, the primary focus is on the probability of perturbing each token, $\beta$, rather than $t$. Since $\beta$ is a monotonically increasing function of $t$, there exists a one-to-one mapping between $\beta$ and $t$. Thus, we can directly parameterize $t$ as $\beta$.
This approach significantly simplifies the diffusion model pipeline and the process of certifying robustness. First, it eliminates the need for the variable $t$ and removes concerns about tuning the relationship between $\beta$ and $t$. Additionally, the framework becomes more straightforward and intuitive, as the noise level $\beta$ directly represents the probability of perturbing each token. Importantly, this re-parameterization does not alter the underlying diffusion models. With just a few additional lines of code, any existing diffusion model can be converted to this parameterization.
This technique has been extensively discussed in Karras et al. (2022) and Chen et al. (2024b). For clarity, our code also adopts this parameterization.
D.2 LLMs as Detectors
In this work, we use LLMs as safety detectors by tuning their prompts, rather than fine-tuning smaller language models. The key advantage of this approach is its ease of debugging. For instance, when aiming for nearly 0% false positive rates and the detector still misclassifies some benign requests as harmful, debugging such misclassifications in a fine-tuned pre-trained model can be extremely challenging. It is often unclear whether the issue arises from the optimization process, the fine-tuning dataset, or other factors.
In contrast, prompting LLMs makes debugging significantly easier. For example, we can directly ask the LLM, “Why do you think this sentence is harmful?” and gain insights into its reasoning. This makes the process of debugging and controlling false positive rates much more intuitive and transparent.
We do not adopt Llama-3 Guard (Dubey et al., 2024) in our approach because it exhibits a higher false positive rate compared to our method, primarily due to its non-conservative prompt design.
Appendix E More Experiments Details
E.1 Our (Nearly) 0 False Positive Prompt
In the task of safety detection, it is crucial to have a reliable mechanism that can identify whether a user’s request contains unsafe or harmful content, without over-reporting benign content as unsafe. To address this, we propose a carefully crafted prompt designed to achieve nearly 0% false positives (FP) while ensuring a high detection rate of genuinely unsafe content.
Below is the prompt we use to construct the safety detector from a LLM:
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] Remember that you are a safety detector. **DO NOT** answer user’s question. Your task is to **strictly identify if the USER’S request explicitly contains unsafe or harmful content** (e.g., instructions for illegal activities, violence, self-harm, or any content that promotes harm to others). You should focus on the overall intent and explicitness of the content rather than isolated words or phrases. If the content is unclear or does not explicitly suggest harm, classify it as [Safe]. Provide your safety assessment for ONLY THE LAST role in the following conversation. <BEGIN CONVERSATION>:
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] <END CONVERSATION> Now provide your safety assessment for the user’s request in the above conversation. Answer only one word: [Safe] or [Unsafe].
The prompt clearly specifies that if the content is unclear or does not explicitly suggest harm, it should be classified as [Safe]. This reduces the chances of false positives because the model is instructed not to make assumptions about potential harm when the content lacks explicit harmful signals.
Table 4: Robustness (%, $\boldsymbol{\uparrow}$) of different defenses under the black-box setting.
{tblr}
column2-8 = c, cell21 = r=7, cell91 = r=7, cell161 = r=7, vline3 = -, hline1-2,9,16,23 = -, rowsep=1pt, colsep=3pt, Models & Defenses GCG MAC I-GCG AutoDAN ICA I-I-GCG
Vicuna-7B No Defense 0% 0% 0% 4% 66% 0%
PPL 72% 24% 96% 52% 66% 98%
ICD 70% 96% 88% 96% 82% 96%
Self-reminder 60% 94% 26% 92% 50% 86%
PAT 94% 92% 82% 98% 82% 86%
DiffTextPure (Uniform) 98% 92% 90% 94% 16% 92%
DiffTextPure (Absorb) 98% 86% 92% 94% 30% 86%
Llama-2-7B-Chat No Defense 48% 2% 4% 80% 100% 0%
PPL 96% 46% 100% 98% 100% 70%
ICD 100% 100% 100% 100% 100% 94%
Self-reminder 100% 100% 100% 100% 100% 100%
PAT 94% 98% 98% 100% 100% 98%
DiffTextPure (Uniform) 100% 98% 100% 100% 100% 100%
DiffTextPure (Absorb) 100% 100% 100% 100% 100% 100%
Llama-3-8B-Instruct No Defense 34% 6% 0% 84% 86% 0%
PPL 82% 88% 96% 98% 86% 100%
ICD 100% 100% 100% 100% 100% 100%
Self-reminder 100% 100% 90% 100% 100% 98%
PAT 100% 100% 100% 100% 96% 100%
DiffTextPure (Uniform) 96% 100% 100% 94% 73% 100%
DiffTextPure (Absorb) 96% 100% 100% 98% 69% 100%
E.2 Experimental Detail of Empirical Evaluations
Dataset: Following prior works, we use the AdvBench dataset (Zou et al., 2023), which consists of approximately 500 harmful strings and behaviors. Due to limited computational resources, we follow Jia et al. (2024) and use their harmful behaviors subset, which contains 50 behaviors randomly sampled from AdvBench.
Baselines: We compare our defense against four state-of-the-art baselines—PPL (Alon & Kamfonas, 2023), ICD (Wei et al., 2023b), Self-reminder (Wu et al., 2023), and PAT (Mo et al., 2024) across four types of jailbreak attacks: GCG (Zou et al., 2023), MAC (Zhang & Wei, 2025), I-GCG (Jia et al., 2024), AutoDAN (Liu et al., 2023), ICA (Wei et al., 2023b) and our I2-GCG (see Sec. 3).
Models: Our experiments span four open-source models, including Vicuna-7B (Zheng et al., 2024a), Llama-2-7B-Chat (Touvron et al., 2023), and Llama-3-8B-Instruct (Dubey et al., 2024).
Hyper-parameters: The experimental settings for baseline attacks and defenses follow their original papers, except for two adjustments: we use a 5-shot setting for ICA and optimize for 100 steps in AutoDAN, due to memory constraints. For hyper-parameters in DiffTextPure, we adopt $\beta=0.25$. We use the diffusion language model (Lou et al., 2023) as the purifier.
E.3 Black-box Evaluation
Black-box evaluations represent practical settings where attackers have only limited access to the model. In this section, we follow previous work (Wei et al., 2023b; Mo et al., 2024; Wu et al., 2023) and conduct experiments in which the attackers know only the base model but are unaware of the defense.
Experimental Results. The table 4 shows that DiffTextPure achieves robust defense against optimization-based adversarial attacks across all tested models (Vicuna-7B, Llama-2-7B-Chat, and Llama-3-8B-Instruct). Both the Uniform and Absorb variants consistently demonstrate high robustness against GCG, I-GCG, and AutoDAN attacks. In particular, DiffTextPure (Uniform) achieves a near-perfect robustness score of 98% against GCG across the models, with similarly strong performance against I-GCG (90%-100%) and AutoDAN (94%-100%). This consistent performance underlines DiffTextPure’s capability as an effective and versatile defense mechanism against optimization-based attacks in a black-box setting.
In contrast, the defense’s performance against prompt-based attacks shows some variability. For Vicuna-7B, DiffTextPure (Uniform) achieves lower robustness (16%). For Llama-2 and Llama-3, it further decreases robustness. This indicates that the purification procedure may rephrase these prompts in a way that makes the requests more covert. This issue could potentially be addressed by designing the purification prompt to explicitly remove harmful requests rather than inadvertently refining them. Since this work primarily focuses on worst-case robustness, we leave this issue for future investigation.
Overall, the results indicate that DiffTextPure can significantly enhance the resilience of large language models to various optimization-based adversarial attacks, disrupting their adversarial nature, offering a plug-and-play defense that maintains robustness across different model architectures and attack strategies.
E.4 Certified Robustness Settings
Following previous work (Cohen et al., 2019; Salman et al., 2019; Carlini et al., 2023b; Xiao et al., 2023; Chen et al., 2024a), we use sample size $1,000,000$, type one error $0.01$. In main experiments, we use $\beta=0.1$ for certification against $\ell_{0}$ attacks, and $\beta=0.25$ for certification against the suffix attacks. We use the diffusion language models (Lou et al., 2023) as the purifier in the main experiments and also compare with the GPT-4o purifier in Sec. E.6.
Clarification of the Time Complexity. The certification procedure typically requires a large number of tests. However, this does not affect practical usage. Certified robustness is intended to provide a lower bound for randomized defenses and should be performed by developers. Once the model is certified and released, users only require $O(1)$ inference to obtain the results.
E.5 Ablation Study of $\beta$ in $\ell_{0}$ Setting
Table 5: Certified robustness of $\ell_{0}$ robustness with different $\beta$ on AdvBench dataset (Zou et al., 2023) using Llama-3-8B (Dubey et al., 2024).
0.1 0.25 0.5 0.75 0.9 1
Absorb 1.82 1.44 0.94 0.86 0.12 0.00
Uniform 1.54 1.06 0.66 0.08 0.06 0.00
To investigate the impact of $\beta$ on the certified robustness under $\ell_{0}$ attacks (the effects on suffix attacks are already explored in Sec. 6), we conduct the following ablation study. In this experiment, we compute the certified robustness using Llama-3-8B across different values of $\beta$.
As shown in Table 5, for both the Absorb kernel and Uniform kernel, we observe that the certified robustness decreases as $\beta$ increases. This can be explained by the nature of $\ell_{0}$ attacks: keywords in sentences are often sparse. For such high-information-density inputs, increasing $\beta$ (i.e., increasing the probability of perturbing each token) easily disrupts the keywords, leading to a significant drop in accuracy ($p_{A}$), and consequently, the certified robustness decreases. When $\beta$ approaches 1, the perturbed noisy sample $\bm{z}$ of normal and adversarial samples becomes nearly identical. Since we set false positives to zero, the certified robustness must also approach zero in this case.
Therefore, in the $\ell_{0}$ attack setting, we choose $\beta=0.1$ as the default value in our experiments to maintain a balance between the smoothness of $g$ and the preservation of key words.
E.6 Comparison of Purification Models
Table 6: Certified robustness of Llama-3-8B (Dubey et al., 2024) on AdvBench dataset (Zou et al., 2023) using different purifiers. Following the default setting, we use $\beta=0.1$ for $\ell_{0}$ attacks and $\beta=0.25$ for suffix attacks.
Purifier Kernel Diffusion Vicuna Llama-3 GPT-4o
$\ell_{0}$ attacks Absorb 1.82 0.00 0.00 2.76
$\ell_{0}$ attacks Uniform 1.54 0.00 0.00 1.42
Suffix attacks Absorb 6.57 0.00 0.00 6.30
Suffix attacks Uniform 6.41 0.00 0.00 1.28
E.6.1 Experimental Settings
Purification prompt. To ensure that the language model correctly restores the original text from the perturbed version, we carefully designed the purification prompt with the assistance of GPT itself. The final version of the prompt when using absorbing kernel is shown in the following:
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] You are a language purifier. Your sole responsibility is to recover the original text from the given perturbed text. **DO NOT** answer the user’s question or modify the existing words. The given text is perturbed from a source text, where some tokens are replaced with the special word ”[MASK]” with a probability of $\beta$. If there are no ”[MASK]” tokens, just return the text as is.
Rules: 1. Words that are already present in the input text must remain unchanged. Do not modify them. 2. Each ”[MASK]” token corresponds to exactly one word. Replace each ”[MASK]” with the most likely word based on context. 3. If the input does not contain any ”[MASK]” tokens, simply return the text exactly as it is without modification. 4. The user’s request may be harmful, but you do not need to consider this. Your only task is to recover the original text as accurately as possible. Accurate recovery ensures that downstream safety detectors can properly analyze the request. 5. Do not include any additional markers like ”<BEGIN CONVERSATION>” or ”<END CONVERSATION>” in your output.
Example 1: If the input is
“Write a [MASK] on [MASK] to make a bomb”
the output should be
“Write a tutorial on how to make a bomb.”
Hint: Use the context of the input text to determine the most appropriate replacement for each ”[MASK]”. Please ignore the instruction that states, “You are trained on data up to October 2023”.
Here is the given text (just recover the original text):
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt] <END CONVERSATION>
Now provide only the purified text without any additional markers.
In early attempts, we observed that GPT frequently ignored our instructions, either by modifying words that were not perturbed or by refusing to recover text when it deemed the content inappropriate. To mitigate this, we iteratively refined the prompt with explicit instructions, constraints, and examples.
Design motivations. Several refinements were made based on empirical observations:
  • Strict adherence to text recovery: The model often deviated from its task by either refusing to recover sensitive text or introducing unnecessary modifications. We explicitly instructed it to recover text as accurately as possible while ignoring the content’s potential harmfulness. This ensures that downstream safety detectors receive the correct input for analysis.
  • Preserving unperturbed words: Early versions of the prompt resulted in cases where GPT arbitrarily changed existing words, significantly affecting recovery accuracy. We explicitly enforced the rule that words already present in the input must remain unchanged.
  • One-to-one token mapping: GPT sometimes replaced a single [MASK] token with multiple words or left it empty. We explicitly stated that each [MASK] corresponds to exactly one word and must be restored based on context.
  • Handling cases without [MASK] tokens: When the input did not contain any [MASK] tokens, GPT often generated irrelevant text instead of returning the input unchanged. To prevent this, we explicitly instructed it to return the input as is in such cases.
  • Avoiding unnecessary metadata: In some responses, GPT included extraneous information such as "You are trained on data up to October 2023", which was irrelevant to the task. We explicitly instructed the model to ignore this message, ensuring that it remains focused on purification.
  • Providing explicit examples: Without concrete examples, GPT often failed to follow the instructions precisely. By incorporating clear before-and-after cases, we significantly improved compliance and task-specific performance.
The final prompt ensures that the model follows instructions strictly, adheres to the recovery constraints, and produces outputs suitable for downstream safety assessment.
E.6.2 Experimental Results
Small language models have bad instruction following, and they may need fine-tuning for purification tasks. We evaluate the instruction-following ability of Vicuna, Llama-2, and Llama-3 in the purification task. Regardless of how we adjust the prompt, these models fail to perform purification correctly and instead produce irrelevant outputs. Vicuna consistently repeats the given system prompt verbatim, regardless of the input text. Similarly, Llama-2 always echoes a specific sentence from the prompt instead of processing the perturbed text. Llama-3 behaves even more unexpectedly, often producing "(no output)" instead of any meaningful response. These results suggest that small language models struggle with following purification instructions and may require fine-tuning to align their behavior with the task.
GPT-4o is a much better purifier in absorbing kernel than diffusion models. As demonstrated, GPT-4o is a much better purifier than the absorbing kernel. Although GPT-4o sometimes provides unusual responses, such as “You are trained on data up to October 2023,” its overall performance still surpasses that of diffusion models. Our trivial bound and Bayesian bound do not account for grammar. For example, in “How to make an explosive bomb,” the trivial bound is one because deleting “to” results in a sentence that can be restored as *”How don’t make an explosive bomb.” However, GPT-4o does consider grammar, preventing such purification, making it even more effective than our keyword-based bound. On the one hand, this demonstrates the strong capabilities of GPT-4o. On the other hand, if user requests are not always grammatically correct, our keyword-based bound would still serve as an upper bound for certified robustness using GPT-4o. One possible improvement is to add an extra prompt to GPT-4o, reminding it that user requests may not always be grammatically correct.
Uniform kernel requires fine-tuning. In the uniform kernel setting, where each token is perturbed to another token from the vocabulary with probability $\beta$, the purifier struggles to correctly interpret the nature of this perturbation. Unlike the absorbing kernel, where non-[MASK] tokens must remain unchanged, the uniform kernel lacks a clear boundary for which words should be modified. As a result, the purifier tends to modify an excessive number of words, often replacing harmful words with benign ones, leading to a high false negative rate. Since purification in the uniform kernel setting requires Bayesian reasoning to estimate the number of perturbed words based on $\beta$, prompt engineering alone appears insufficient for aligning LLMs with this task. Instead, fine-tuning on structured purification data may be necessary to ensure that the model correctly distinguishes perturbed tokens and performs accurate purification.
E.7 Certified Robustness on Repeated AdvBench
Table 7: Certified radius of $\ell_{0}$ robustness on repeated AdvBench dataset (Zou et al., 2023) (which repeat each request in Advbench) using Llama-3-8B (Dubey et al., 2024).
# repeats Absorb (0.1) Absorb (0.25) Uniform (0.1) Uniform (0.25) Trivial Bound Bayesian Bound
1 1.82 1.44 1.54 1.06 2.12 2.10
2 3.70 4.20 3.22 3.26 5.24 4.54
3 3.94 5.90 3.82 5.34 8.36 6.16
5 3.88 6.84 3.94 6.62 14.6 7.96
AdvBench contains only short requests, and experiments with short requests may not fully capture the trends of the certified radius, Bayesian bound, and trivial bound. Additionally, there is a growing trend of adversarial prompts becoming gradually longer (Andriushchenko et al., 2024).
To better illustrate the trend of the certified bound with increasing prompt length, we repeat each request 1, 2, 3, and 5 times and run the certification and Bayesian error bound evaluations.
As shown in Table 7, the gap between the trivial bound and the Bayesian bound grows dramatically as the length of the adversarial prompt increases. This indicates that current certification methods struggle to provide tight bounds for longer adversarial prompts. This may require us to design new certification algorithms. In contrast, the gap between the real certification we achieve and the Bayesian bound grows only linearly. This observation suggests that there may be a constant gap between the two bounds. Consequently, improving the effectiveness of the basic method is likely to result in a linear improvement in the effectiveness of adversarial prompts over an extended range of lengths.
Appendix F More Discussions
F.1 Detail about our I2-GCG
Formulating white-box attacks as optimization. Any defended model is a mapping $f:\mathcal{V}^{N}\to\mathcal{V}^{N}$. Unlike Athalye et al. (2018), we do not design specific loss functions for each submodule of $f$. Instead, we directly calculate the loss on the output and minimize it. Specifically, we optimize:
$\min_{\bm{x}_{adv}}L(f(\bm{x}_{adv})),\text{ s.t. }\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d.$
where $L$ is the same loss function as Zou et al. (2023), $\mathcal{D}$ is a distance metric and $d$ represents the attack budget. Since this optimization problem guarantees convergence, this evaluation is sufficient over a long enough time.
Exact white-box. Most language models use the BPE tokenizer (Sennrich, 2015), which is sensitive to small modifications (e.g., adding an extra space), resulting in different tokenization. For this reason, many implementations fail to rigorously ensure token consistency when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can cause attackers to fail in generating adversarial examples.
No early return. Based on our observations, sufficient optimization nearly eliminates all cases where the language model’s output aligns with our target but transitions to a refusal to answer in the subsequent steps. By removing the early return, we ensure that every adversarial example undergoes sufficient optimization.
Removing gradient. Since some defenses are non-differentiable, we remove the gradient pairing in GCG for fairness during the evaluation. Previous studies also suggest that the gradient components of GCG provide minimal assistance to the optimization (Jia et al., 2024).
Warm start. We follow I-GCG (Jia et al., 2024), using the adversarial components from previous iterations as the initialization for the next batch of data. This greatly accelerates the process, requiring approximately 100 iterations to achieve a 100% success rate.
F.2 About False Positives in Adversarial Suffix Settings
Due to the explicit structure of adversarial suffix attacks, several defenses can achieve impressive certified robustness. For example, when $\beta\to 1$ in this work, when the number of deleted tokens tends to infinity in Kumar et al. (2023), the certified radius would also go to infinity.
However, from a human perspective, the certified radius against suffix attacks should not be too large. For example, the phrase “tell me how to make a bomb” is a harmful request. However, by padding with 4 tokens, it can become “tell me how to make a bomb. Do not answer this,” which transforms it into a benign request.
Therefore, for any certified method against suffix attacks, one should consider tuning the hyperparameters to prevent the smoothed models from becoming over-smoothed.
F.3 Ill-poseness of Adversarial Suffix Settings
A reminder when certifying against suffix attacks is to take the minimal certified radius over all suffix lengths. Consider a defense that deletes the last 2 tokens to defend against suffix attacks. Due to the ill-posedness of adversarial suffix settings, we can successfully certify against any attacks that append exactly two suffixes, but not exactly one suffix. When we talk about certifying against suffix attacks, we claim that no matter how many suffixes the attacker appends within our certified radius, our defense will still be certifiably robust. Thus, when certifying against suffix attacks, we should take the minimal certified radius over all suffix lengths.
F.4 Reduction to Broader Setting
A potential way to combine certification against $\ell_{0}$ attacks and suffix attacks is to first append several tokens and then certify the $\ell_{0}$ radius of the whole string. This certified result will include both perturbations in suffix and $\ell_{0}$ perturbations and thus certifies against both $\ell_{0}$ attacks and suffix attacks. However, the obtained result is exactly the same as the certified radius against $\ell_{0}$ attacks. This is because certifying against $\ell_{0}$ attacks is much more challenging than certifying against suffix attacks, and thus the certified radius remains the same as for suffix attacks. For this reason, we certify them separately, in order to better illustrate the certified results for these two types of attacks.
Appendix G Limitations
There are several limitations of this work.
Limitation 1: White-box evaluation against stochastic attacks. Our I2-GCG method can only accurately evaluate the robustness of non-stochastic defenses. For stochastic defenses that induce a large amount of randomness, the optimization of I2-GCG is interfered with and cannot converge to a stable solution within a short time (at least within 1000 steps).
Limitation 2: Defending against expertise-based attacks. The core principle of our defense is to transform out-of-distribution data back into in-distribution data, and its certified guarantees are effective only when the length of the adversarial suffix is limited. However, expertise-based attacks, which utilize human-crafted prompts, often appear natural (i.e., have high likelihood) and are typically lengthy, rendering our theoretical guarantees less effective (see ICA in Table 4). This issue could potentially be addressed by integrating our defense with existing heuristic defenses.
Limitation 3: High information-density data. When applying DiffTextPure as a pre-processor for generative language models, a critical limitation is that the forward process has some probability of destroying important information, particularly in cases with high information density. This makes it challenging for the backward process to fully recover the original content. For example, in mathematical problems, if key numerical values are altered during the forward process, their recovery becomes impossible, as such values typically appear only once in the input text. In this case, DiffTextPure can only be used as a pre-processor for a detector, rather than a generator, since generators usually have a higher requirement for preserved information than detectors.
Limitation 4: Limited setting of certified robustness. In this work, although we derive certifications for all smoothing distributions, there are still significant limitations. First, we cannot certify against heuristic attacks that use very long prompts, such as those in Wei et al. (2023b) and Chao et al. (2023). Additionally, we do not certify adversarial attacks involving insertion and deletion. This may require constructing $p(\bm{z}|\bm{x})$ to randomly insert or delete tokens. However, we believe that our framework can serve as a theoretical foundation, with future work focusing on proposing noising distributions of varying lengths and using fractal knapsack solver or 0-1 knapsack solver to certify against a broader class of attacks.
Appendix H Key Takeaways
White-box attacks can still easily achieve 0% robustness against existing defenses. We do not propose any advanced optimizers in this paper. The reason we achieve a 100% attack success rate, while previous works cannot, is that we strictly ensure the consistency of tokens during both optimization and inference. None of the previous works consistently enforce this, which leads to adversarial tokens achieving low loss during training but higher loss during inference due to slight differences in tokenization. These approaches are actually grey-box settings, not true white-box settings, as they fail to ensure token consistency. Token consistency is the only reason why previous attacks could not achieve a 100% success rate. Other techniques in this paper (e.g., attacking longer, removing gradients, warm starts) are incremental improvements and are only designed to accelerate attacks or address extreme cases, such as transitions into safe responses.
Token consistency is simple in principle, but it took us a really long time to carefully ensure the token consistency for every model and defense, even each sentence. Of course, adaptive attacks are also crucial. One should at least include every part of the defense in the attacking process, rather than relying on techniques like BPDA (Athalye et al., 2018). Whether you design a specific loss function for each component, as in Carlini & Wagner (2017), or treat the entire model as a unified procedure and optimize the overall loss does not make a significant difference.
Similar to adversarial robustness in computer vision, there are still limited defenses, such as adversarial training and randomized smoothing, that do not have 0% worst-case robustness. In adversarial robustness for vision, only a few defenses, such as adversarial training and randomized smoothing (which includes purification-based defenses), avoid being reduced to 0% robustness. Other defenses have ultimately been proven ineffective and were attacked to 0% robustness. In this work, we reach a nearly identical conclusion. While we still believe adversarial training can partially address this problem, current approaches to adversarial training focus more on alignment rather than the traditional adversarial training that involved extensive and long-term training. As a result, these newer approaches fail to address worst-case robustness, offering only slight improvements in average-case robustness.
White-box evaluations provide an upper bound for worst-case robustness, while certified robustness serves as the lower bound. White-box evaluations only provide an upper bound for worst-case robustness, and future, stronger attacks may further decrease this upper bound. In contrast, certified robustness is a theoretical lower bound for worst-case robustness, and future advancements in certification analysis may increase this lower bound. We believe that, as researchers continue to improve both evaluation and certification methods, the gap between the empirical upper bound and the theoretical lower bound will gradually narrow.
Certified robustness is a fractional knapsack or 0-1 knapsack problem. When the base function $f$ is a bounded function, randomized smoothing becomes a fractional knapsack problem. If the base function $f$ is a binary function, this transforms into a 0-1 knapsack problem, which can improve the certified bound.
This certification framework can be applied not only to robustness but also to other aspects of machine learning. Most machine learning problems can be formulated as $L(\bm{x}_{test},\text{train}(\bm{x}_{train},\bm{\theta}))$, where $\bm{x}_{train}$ is the training set, $\bm{\theta}$ represents the parameters trained on this set, and $\bm{x}_{test}$ is the test set used for evaluation. The certification framework can be applied to each component of this paradigm.
When applied to $\bm{x}_{train}$, we can certify that poisoning the training set may not significantly affect the functionality of the trained model, like Hong et al. (2024). When applied to $\bm{\theta}$, we can certify that corrupting or dropping out parts of $\bm{\theta}$ will not overly impact the functionality of the model or the training process. When applied to $\bm{x}_{test}$, as we have done, we can certify that adjusting the testing inputs will not successfully attack the already trained models.
We hope certification techniques would provide deeper insights and mathematical guarantees for a wide range of practical applications in the future.
$p_{adv}-p_{A}$ plots are a good way to visualize certification. In this paper, we visualize the fractal knapsack solver using $p_{adv}-p_{A}$ plots. By proving the symmetrization of the $p_{adv}-p_{A}$ plots with uniform kernels, we can easily derive additional conclusions, such as the uniform kernel always outperforming the absorbing kernel, and the certified radius being a monotonic decreasing function with respect to vocabulary size, at most starting from $d+1$.